### v1

# Towards Context-Aware Domain Generalization: Representing Environments with Permutation-Invariant Networks

Abstract
In this work, we show that information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models when applied in new domains or production environments. We formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same environment/domain as the input itself. These representations are jointly learned with a standard supervised learning objective, providing incremental information about the unknown outcome. Furthermore, we offer a theoretical analysis of the conditions under which our approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice. Additionally, we contribute insights into the kind of distribution shifts for which our approach promises robustness. Our empirical evaluation demonstrates the effectiveness of our approach for both low-dimensional and high-dimensional data sets. Finally, we demonstrate that we can reliably detect scenarios where a model is tasked with unwarranted extrapolation in out-of-distribution (OOD) domains, identifying potential failure cases. Consequently, we showcase a method to select between the most predictive and the most robust model, circumventing the well-known trade-off between predictive performance and robustness.
1 Introduction
Figure 1: Conceptual sketch of our setup and approach. A) Data-generating process (DGP) that fulfills our criteria. We assume that the environment $E$ is a source node that is not caused by any system variable and that the relationship between $\mathbf{X}$ and $\mathbf{Y}$ varies with the environment. $\mathbf{S}^{(n)}$ is a set of $n$ IID inputs available in the new environment (i.e. context). B) The source component shift corresponding to the assumed DGP and example data coinciding with Simpson’s paradox. C) The workings of our approach in a test environment. A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context. An inference network (e.g., a classifier) processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome (e.g., label) of the target input. The set-representation can be combined with the input to reliably detect out-of-distribution (OOD) queries and prevent failure cases in domain generalization due to model misspecification.
Distribution shifts are the cause of many failure cases in machine learning [27, 32] and the root of various peculiar phenomena in classical statistics, such as Simpsons’ paradox [45, 57]. In this work, we employ permutation-invariant neural networks as set-encoders [15, 7] to improve the predictions of standard supervised models under distribution shift. Specifically, we consider training in the realm of Domain Generalization (DG), a setting where data from distinct environments [^1] We use the terms environment and domain interchangeably. is available for training and testing [39, 74].
For illustration, consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$. Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands. It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$. This raises two questions:
  • (1)
    Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$?
  • (2)
    How should $E$ be represented to maximize the performance gain?
The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$. For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region. Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically.
A straightforward answer to the second question is to distinguish environments by discrete labels. However, we argue that $E$ should be a continuous embedding. First, continuous embeddings can also be computed for new environments that have not been present in the training data. Second, when $p(\mathbf{Y}\mid\mathbf{X},E)$ receives a continuous $E$, it can potentially configure itself for unseen environments by interpolating between the training environments. And finally, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
In the current work, we systematically investigate both questions, formalize three criteria when $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and demonstrate how continuous embeddings $E$ can be learned from auxiliary data by means of set encoders (see Figure 1). Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach. First, we require that a single input alone is insufficient to deduce the originating environment (see Criterion 2.2). If this condition is not met, our model cannot possibly extract more information about the originating environment, compared to a standard model. Second, we require that, if the environment label is known by an oracle and given as an additional input to the standard model, its performance must improve (see Criterion 2.3). Otherwise, the set-based representation of the environment cannot possibly yield benefits. Notably, these two criteria are easy to verify and also met under the source component shift (see Section 2.6), which is a dataset shift that occurs in many scenarios [46].
When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and the potential to generate silent failures. While our approach is not exempt from this “curse of extrapolation”, it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [41]. Moreover, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs. models that are robust to out-of-distribution (OOD) scenarios on the fly. Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [65, 42, 37]. Accordingly, we can adaptively select the most robust model in the OOD setting and the most predictive model in the ID setting, an approach we demonstrate on the ColoredMNIST data set [4]. In summary, our contributions are:
  • We propose a novel approach to Domain Generalization (DG) that leverages context information from new environments in the form of learnable set-representations;
  • We formalize the necessary and empirically verifiable conditions under which our approach can reap benefits from context information and improve on standard approaches;
  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required.
2 Method
2.1 Notation
We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively. We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ an i.i.d. sample (i.e., a set of further inputs) from the given domain. The domain label $E$ is only known during training time and unknown during inference.
2.2 General Idea
The key idea of our approach is to build models that utilize not only a singleton input $\mathbf{X}$ to predict a target $\mathbf{Y}$, but also information about the environment of $\mathbf{X}$ that can improve the prediction $\hat{\mathbf{Y}}$. Providing environmental information in the form of a one-hot label is hardly feasible, as it presupposes the exact number of possible environments to be known during training, and that we always know from which environment an input originates at inference time. Consequently, such an approach is doomed to fail when the test data originates from a novel or unknown environment.
To overcome this problem, we employ permutation-invariant neural networks to adaptively represent environmental information given a set of test inputs. We will first detail our approach and then discuss criteria under which we can expect to reap benefits from the additional set-representation. Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which our approach may prove advantageous in practice. Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
2.3 Permutation-Invariant Neural Networks
As mentioned above, a basic goal of our approach is to synthesize contextual information about a target input $\mathbf{X}$ by compressing a set of $n$ further inputs $\mathbf{S}^{(n)}:=\{\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n}\}$ from the same environment into a permutation-invariant representation. The notion of permutation invariance is closely related to a core concept in probabilistic modeling and Bayesian inference – exchangeability [44]. Accordingly, an exchangeable sequence of random vectors is characterized by a joint distribution which is invariant to any permutation of the elements:
$p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n})=p(\mathbf{X}_{\pi(1)},\mathbf{X}_{\pi(2)},\dots,\mathbf{X}_{\pi(n)}),$ (1)
where $\pi:\mathbb{N}\rightarrow\mathbb{N}$ denotes an arbitrary permutation of index elements $n\in\mathbb{N}$.
Exchangeable observations may come in various forms, for instance, patients entering a hospital, different measurements obtained with the same device, or sets of visual images, such as faces in a crowd. When it comes to learning exchangeable symmetries, permutation-invariant functions can serve as a key building block in neural architectures [7], as they automatically encode a favorable inductive bias towards permutation-invariance by design. A simple and intuitive way to create permutation-invariant functions involves the sum-decomposition
$h(\mathbf{S}^{(n)})=\rho\left(\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})\right),$ (2)
where $\sigma$ and $\rho$ can be any functions, including deep neural networks [69, 7]. $h$ is permutation-invariant due to the summation operator which ensures that the argument of $\rho$ is agnostic to the order of elements in the set $\mathbf{S}^{(n)}$.
Despite having favorable theoretical properties, plain sum-decompositions can have limited representational capacity in practice [69, 58]. Thus, more expressive permutation-invariant functions can be learned by stacking equivariant transformations with sum-decompositions [69, 7] or by using a self-attention mechanism without positional encodings [33]. In order to re-assess the expressiveness of these methods, we show a comparison of the binary domain classification accuracy based on domain overlap and set size in Figure 10.
2.4 Context-Aware Model
Our model consists of two key components (also illustrated in Figure 1):
  • a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and
  • an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction.
The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short. For a given supervised learning task, we aim to find the minimum to the following optimization problem
$\boldsymbol{\theta}^{\star}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{p(\mathbf{X},\mathbf{Y},E)}\left[c(f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)}),\mathbf{Y})\right],$ (3)
where $c$ is a task-specific loss function (e.g., cross-entropy for classification or mean squared error for regression). The algorithmic details for optimizing Equation 3 are detailed in Algorithm 1. For practical reasons, we first apply a feature extractor $g$ and then pass $\{g(\mathbf{X})\mid\mathbf{X}\in\mathbf{S}^{(n)}\}$ as input to the set encoder and $g(\mathbf{X})$ as input to the inference network, building upon the features extracted by $g$. Note, that $g$ can be a pre-trained network, in which case we can treat it as a fixed transformation, or its parameters can be optimized jointly with $\boldsymbol{\theta}$.
Data: Samples from the joint distribution $p(\mathbf{X},\mathbf{Y},E)$
Input : Composite model parameters $\boldsymbol{\theta}$, set size $n$, batch size $m$, loss-function $c$, number of iterations $k$, learning rate schedule $\alpha(k)$
1 for $i=1,\dots,k$ do
2       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
3       for $j=1,\dots,m$ do
4             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
5             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
6       end for
7      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
$\boldsymbol{\theta}_{k}\leftarrow\boldsymbol{\theta}_{k-1}-\alpha(k)\nabla_{\boldsymbol{\theta}}\left(\sum_{j=1}^{m}c\left(f_{\boldsymbol{\theta}}(\mathbf{x}_{j},\mathbf{s}_{j}^{(n)}),\boldsymbol{y}_{j}\right)\right)$
8 end for
Output : Trained context-aware model $f_{\boldsymbol{\theta}}$
Algorithm 1 Optimizing Equation 3 for context-aware domain generalization.
2.5 Criteria for Improvement
In the following, we establish criteria under which our method can exploit the distribution shifts between environments and yield improved predictions. In total, we propose three criteria that are necessary to achieve incremental improvement. In Theorem 2.1, we show how these criteria are related to each other. In the formulations below, $I(\mathbf{X};\mathbf{Y})$ denotes the mutual information between random vectors $\mathbf{X}$ and $\mathbf{Y}$ and $I(\mathbf{X};\mathbf{Y}\mid\mathbf{Z})$ denotes the conditional mutual information given a third random vector $\mathbf{Z}$. The symbol $\perp$ (resp. ⟂$\not\perp$̸) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp. dependent) or conditionally independent (resp. dependent) given a third random vector $\mathbf{Z}$.
First, we require that given an input $\mathbf{X}$, a further set of i.i.d. inputs $\mathbf{S}^{(n)}$ from the same environment provides incremental information about $\mathbf{Y}$. This is exactly what we need to achieve improved predictive performance, and we can formally define it as our first criterion:
Criterion 2.1.
$\mathbf{S}^{(n)}\not\perp\mathbf{Y}\mid\mathbf{X}$ or equivalently $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})>0$.
The second criterion requires that, given a target input $\mathbf{X}$, a set of further i.i.d. inputs $\mathbf{S}^{(n)}$ from the same environment provides additional information about the origin environment of $\mathbf{X}$.
Criterion 2.2.
$E\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ or equivalently $I(E;\,\mathbf{S}^{(n)}\mid\mathbf{X})>0$.
In Figure 1, an instance $\mathbf{X}$ cannot be assigned with complete certainty to an environment. Consequentially, further data provides additional information about the environment. In general, the more data we consider, the better we can predict the originating environment. Crucially, this criterion is not satisfied, if we can recover the origin environment from the singleton input $\mathbf{X}$ alone.
The third criterion requires that the singleton input $\mathbf{X}$ gains information about $\mathbf{Y}$ if we also consider the origin environment $E$ of $\mathbf{X}$.
Criterion 2.3.
$\mathbf{Y}\not\perp E\mid\mathbf{X}$ or equivalently $I(\mathbf{Y};\,E\mid\mathbf{X})>0$.
In Figure 1, this is evidently the case: If we knew the environment from which the data stems, we could improve our prediction of $\mathbf{Y}$. Furthermore, this criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy.
In what follows, we show that Criterion 2.2 and Criterion 2.3 are necessary conditions for Criterion 2.1. We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1. We even generalize this result for the case where the environment label is not inferable with 100% accuracy.
Theorem 2.1.
The following statements hold:
  • (a)
    If $E\perp\mathbf{S}^{(n)}\mid\mathbf{X}$, it follows that $\mathbf{Y}\perp\mathbf{S}^{(n)}\mid\mathbf{X}$. This is equivalent to the implication that if Criterion 2.2 is unattainable, then Criterion 2.1 is also not satisfied.
  • (b)
    If $E\perp\mathbf{Y}\mid\mathbf{X}$, we achieve $\mathbf{Y}\perp\mathbf{S}^{(n)}\mid\mathbf{X}$. This statement corresponds to: Criterion 2.3 is a necessary condition for Criterion 2.1.
  • (c)
    Assume that there exists a deterministic function $g$ with $g(\mathbf{S}^{(n)})=E$, then $\mathbf{Y}\not\perp E\mid\mathbf{X}$ implies $\mathbf{Y}\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$. This conveys that if we could perfectly infer $E$ from $\mathbf{S}^{(n)}$, then Criterion 2.3 implies Criterion 2.1.
  • (d)
    Assume that there exists a function $g$ and a noise variable $Z$ that elicits the relation $E=g(\mathbf{S}^{(n)})+Z$ and satisfies $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X}$ as well as $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$. Furthermore, assume that $\mathbf{Y}\not\perp E\mid\mathbf{X}$ and $I(\mathbf{Y};E\mid\mathbf{X})>I(Z;\mathbf{Y}\mid\mathbf{X})$. Then, we achieve $\mathbf{Y}\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$, recovering Criterion 2.1.
The proof for this theorem and a discussion of its presumptions can be found in Section A.2. Unfortunately, we cannot conclude that $Y\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general. An example where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in Section A.1.
2.6 Source Component Shift
Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satifsied. Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [46, Chapter 1.9]. The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$. Problems that conform to the graph in Figure 1 have two important implications. First, the input distribution changes whenever the environment changes. [^2] In our case, we require a stronger assumption, namely, that the domains are not deducible from a single input, as formalized in Criterion 2.2. Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1). For more details on this kind of distribution shift, we refer the reader to [46, Chapter 1.9]. It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [45, 57], which supplies a proof-of-concept for our approach (see Experiment 1).
2.7 Detection of Novel Environments
During test time, data could either originate from an environment that corresponds to one of the training environments (but its origins are unknown) or from a previously unseen environment. In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation. Following [41], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$. As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder. Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment. Following the approach in [41], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment. To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set. We also compare our novel environment detector with the same score function computed solely from singleton features $g(\mathbf{X})$. These results can be quickly previewed in Table 2.
3 Related Work
3.1 Domain Generalization
The aim of domain generalization (DG) is to train models that generalize well under distribution shifts [39, 74]. The DG setting involves access to data from multiple domains during training exploiting the heterogeneity between domains. In this context, the plethora of algorithms aimed at improving robustness has been divided into three categories [60]: (i) algorithms that enhance the data (e.g., via data augmentation or domain synthesis [56, 68]), (ii) algorithms that adapt the learning algorithm (e.g., self-supervised learning [29, 8]), and (iii) algorithms aiming to find robust representations (e.g., via invariance learning [4, 42]). Our approach explores a different avenue exploiting contextual information about the origin of the data via an adaptive environment embedding.
In contrast to domain adaptation (DA) [61], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training. As a middle-ground between DA and DG, test-time adaptation (TTA) involves the provision of unlabeled samples during test time, enabling further fine-tuning of the model [35]. TTA can be categorized into test-time domain adaptation, test-time batch adaptation (TTBA), and online test-time adaptation (OTTA) [35].
Among these settings, our work aligns most closely with the TTBA scenario concerning the utilization of environment information. In TTBA a pre-trained model is adapted to one or a few inputs [54, 72, 51]. Each adaptation depends on the mini-batch at hand. Similarly, we consider a mini-batch as a set input to deliver contextual information. However, we do not adapt or fine-tune our model at test time, but rather directly extract context information via the set-encoder. This allows us to identify whether extrapolation is necessary, enabling model misspecification detection [50]. Moreover, since we do not need to fine-tune the model at test time, our approach is considerably more efficient.
Finally, [12] assume a setting where inputs from multiple domains are available, but it is unknown which sample belongs to which environment – even during training time. The authors infer potential domain labels that are used for downstream invariant learning, for example, via Invariant Risk Minimization (IRM) [4]. This method improves on baseline models and enables the application of DG methods in the absence of domain labels. Interestingly, it could be combined with our approach when no environment labels are provided during training, and we leave this as an avenue for future research.
3.2 Learning Permutation-Invariant Representations
Analyzing set-structured data with neural networks has received much theoretical [58, 7, 43] and empirical [69, 33, 70] momentum in recent years. For instance, [70] build on the set transformer architecture [33] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling. The resulting “PICASO” blocks include consecutive multi-head attention stacks with skip connections, which update their template(s) depending on previous template(s) and the particular input set (in contrast to global templates). Differently, [13] propose to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion. The authors also show how to use their criterion for generative and few-shot classification tasks In a somewhat similar vain, [5] investigate a gradient-based optimization method for aggregating set-structured data. Importantly, these methods comprise a pool of algorithms which can be used as the backbone architecture for realizing the set-encoder in our approach.
Model Symbol Description Purpose
Context-aware model (ours) $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts Y from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Improve predictions
Baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$ Predicts Y from $\mathbf{X}$ alone Verify improvement
Environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ Predicts Y from $\mathbf{X}$ and $E$ Verify Criterion 2.3
Contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $E$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Verify Criterion 2.2
Baseline environment model $f^{E\mid\mathbf{X}}$ Predicts $E$ from $\mathbf{X}$ alone Verify Criterion 2.2.
Table 1: The five different model types used to evaluate our approach and verify the theoretical criteria for improvement.
Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks. In contrast, probabilistic models attempt to learn a conditional or a marginal distribution over the set-representations. In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [44]. In particular, hierarchical or multi-level Bayesian models are used to model the dependencies in nested data, where observations are organized into clusters or levels [64, 21], mirroring the notions of domain or environment in DG. Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
From a somewhat different Bayesian perspective, IID data represents the starting point for learning invariant summary statistics for parameter estimation or model comparison. The pioneering work on Neural Statisticians [14] tackles the task from a variational perspective, learning global set-representations as part of a generative model. Neural processes [19, 30] comprise a related family of set-based models for prediction and uncertainty quantification in supervised learning tasks, rooted in the spirit of Gaussian processes [48]. Finally, [31] explore a permutation-invariant variational autoencoder (SetVAE) with multiple latent variables trained with a modified ELBO objective. The goal of this and follow-up models [71] is accurate generative performance, less so the learning of compact representations. The theoretical and empirical implications of learning random set-representations falls outside the scope of our current work, but offers a potentially fruitful avenue for future research.
Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations. For the purpose of achieving this goal, we employ variants of the DeepSet [69] and SetTransformer architectures for our backbone set-encoder throughout all experiments. Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
3.3 OOD Detection and Selective Classification
Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [3, 67, 52, 23, 66]. Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [67]. OOD detection is closely related to inference with a reject option (also termed selective classification) [20, 16], which allows classifiers to refrain from making a prediction under ambiguous or novel conditions [25]. The reject option has been extensively studied in statistical and machine learning [24, 18, 22, 62], with early work dating back to the 1950s[11, 10, 24].
More recently, [41] explored the utility of selective classification in DG settings. They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent. Post-hoc methods rely on various aspects of pre-trained model outputs, such as the softmax outputs (e.g., [28]), logit outputs (e.g., [26, 36]), or intermediate features (e.g., [2, 53, 59]). In this work, we consider a post-hoc score based on the $k$-nearest neighbors to the training set in feature space similar to [53], which is applicable to both classification and regression settings. Unlike the approach taken in [41], where the focus lies on individual input features, we consider the set summary provided by the set-encoder. Thus, we can identify novel environments even when singleton inputs lack sufficient information.
4 Experiments
Figure 2: Experiment 1. Left: Toy dataset that conforms to our theoretical criteria. A detailed description can be found in Section C.1. Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3. Marginal distributions of the five environments are shown on top. Right: Comparison of environment classification accuracy for different set sizes.
In the following, we explore various aspects of our approach across three different dimensions. First, we show on two datasets that our model achieves improved performance in the ID as well as the OOD setting compared to a baseline model. Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model. We also show that novel environment detection can be utilized to avoid failure cases. Third, we demonstrate that the necessary criteria (see Section 2.5) can be validated empirically, identifying cases where no benefits of our method can be expected. Experimental details can be found in the Appendix.
4.1 Evaluation Approach
To approximate Criterion 2.1, Criterion 2.2 and Criterion 2.3, we are required to train five distinct models (see Table 1 for an overview). We denote our composite model as $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ (see Figure 1) and the baseline model (having no access to the context) as $f^{\mathbf{Y}\mid\mathbf{X}}$. Based on these two models, we can compute the relative improvement achieved by our model relative to a baseline model via
$\mathcal{R}_{\text{I}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (4)
Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$. $\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by our approach and therefore the fulfillment of Criterion 2.1. In the regression setting, we consider the negative L2-Loss as the performance measure.
To validate Criterion 2.2, we train a contextual environment model (referred to as $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$) utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$. Additionally, we train a baseline environment model (denoted as $f^{E\mid\mathbf{X}}$) aimed at predicting $E$ solely from $\mathbf{X}$. We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^3] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix B of the contextual environment predictor relative to the baseline environment predictor. $\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied. In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
Similarly, we consider an environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ that aims to predict $\mathbf{Y}$ from the singleton input and the environment label $E$. We define the relative improvement $\mathcal{R}_{\text{III}}$ of the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ compared to the baseline method $f^{\mathbf{Y}\mid\mathbf{X}}$. In this case, the relative improvement $\mathcal{R}_{\text{III}}$ is associated with Criterion 2.3.
4.2 Experiment 1: Toy Example
Setup
To set the stage, we consider a dataset that is inspired by [63]. The dataset includes data from five environments, defined by five 2D Gaussian distributions differing only in their locations (i.e., mean vectors). The data-generating process is thus:
$j$ $\sim\text{Categorical}(p_{1},\dots,p_{J})$ (5)
$\mathbf{X}^{(ij)}$ $\sim\mathcal{N}(\boldsymbol{\mu}^{(j)},\mathbf{\Sigma})\quad\text{for}\quad i=1,\dots,I_{j}$ (6)
This is the classical setting of Simpsons’ paradox, for which a naive linear model fit oblivious to the hierarchical data-generating process will yield results opposite to the true relationship between $X_{1}^{(j)}$ and $X_{2}^{(j)}$ in each environment (cf. Figure 2 and Section C.1 for details).
We designate the first dimension, $X_{1}$, as the input feature, while the second dimension, $Y=X_{2}$, serves as the outcome variable. Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X_{1}$, obtained by projecting the entire sample onto the $x$-axis in Figure 2. Thus, this setting aligns with Criterion 2.2, and, additionally, corresponds to Criterion 2.3, since learning about the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
Figure 3: Experiment 1. Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on toy example. We also show I (OOD) on OOD data. II depicts the relative improvement of the environment-oracle model compared to the baseline model. III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it. Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
Results
As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input. This evaluation involves studying the classification accuracy in distinguishing between the environments. In Figure 2 we observe that the additional set input improves the ability to predict the environment significantly and the more samples we include, the better the prediction. As expected, a decrease in the distance between environments necessitates more samples to differentiate between environments. Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix G.
Next, we assess the predictive capabilities of our approach across all possible scenarios of “leave-one-environment-out”. This involves training on all environments except one and treating the excluded environment as a novel OOD scenario. Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in Section C.3). We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3. Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3). We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I). This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself.
Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4. Our model captures and utilizes the characteristics of each environment for prediction. In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences. Note that we obtained the best results on this problem by considering a class of linear models that aligns with the data-generating process. However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see Section C.3).
Figure 4: Experiment 1. Predictions performed on the toy dataset illustrated in Figure 2. We show predictions made by both our set-encoder approach and the vanilla model in the ID and OOD settings.
4.3 Experiment 2: ProDAS Example
Setup
We utilize the ProDAS library [40] to generate high-dimensional image data that meets our dataset requirements. The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size. Additionally, the background varies in color and texture, resulting in a complex scenario. We consider the task of predicting the object size. Difficulties arise due to the presence of distinct environments with varying characteristics. Specifically, depending on the environment, a constant is added to the observed object size to get the actual target variable that we aim to predict:
$Y_{\text{gt}}=Y_{\text{observed}}+j\cdot\text{const}_{1}$ (7)
Here, $j\in\{1,2,3,4\}$ denotes the environment, while $Y_{\text{gt}}$ represents the ground truth (or factual) size, obtained as a sum of the observed size $Y_{\text{observed}}$ (relative to the image frame) and a constant depending on $j$. The background color follows a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{j};\boldsymbol{\Sigma})$ where the mean depends on the environment in the following way: $\boldsymbol{\mu}_{j}=\boldsymbol{\mu}_{0}+j\cdot\text{const}_{2}$. Here we assign a small value to $\text{const}_{2}$ to enforce the background distributions to overlap between different environments. Specifically, this construction implies that the relation between input $\mathbf{X}$ and target $Y$ differs across environments. This corresponds to Criterion 2.3. Notably, inferring the originating environment from a single sample is unattainable due to overlapping background distributions (corresponding to Criterion 2.2). Samples of different environments are shown in Appendix D. This example could be inspired by microscopy data where different microscopes correspond to distinct environments, each exhibiting its own characteristics. During training, we assume to have access to the ground truth value $Y_{\text{gt}}$.
Results
In line with the results from the previous toy example, we can demonstrate a strong relative improvement in the ProDAS dataset, as depicted in Figure 5. All formal criteria are satisfied and a very significant improvement is achieved, both in the ID and the OOD setting, by considering the contextual information from the environment. Details for this experiment can be found in Appendix D.
4.4 Experiment 3: Colored MNIST
Setup
The ColoredMNIST dataset [4] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1). Furthermore, there exists label noise, so only in 75% of all cases, the label can correctly be predicted from the input image. To make things more challenging, the image background can take two colors that are also associated with the image label. In the first environment, the association is 90% and in the second one 80%. Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape. However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [42]. For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%). In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
Figure 5: Experiment 2: Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means, no improvement is achieved) on ProDAS dataset. We also show I (OOD) on OOD data. II depicts the relative improvement of the environment-oracle model compared to the baseline model. III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it. Variations arise from using different seeds to partition the ID data into training, test and validation set.
[…]
In-Distribution Out-of-Distribution
Baseline $\mathbf{84.6}\pm 0.3$ $10.2\pm 0.3$
Invariant $72.8\pm 0.9$ $\mathbf{73.1}\pm 0.2$
Selection (Ours) $\mathbf{84.1}\pm 0.3$ $\mathbf{73.1}\pm 0.2$
Selection (Baseline) $\mathbf{84.0}\pm 0.3$ $14.0\pm 0.4$
Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
Table 2: Experiment 3. Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs. The features extracted by our model allow for improved OOD detection compared to the features of the baseline model. Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
Results
Here, we assume the invariant model to be given (see Appendix E for details), but it could also be obtained by invariant learning, e.g. Invariant Risk Minimization [4]. With our novel environment detection approach (see Section 2.7) we can get the best of both worlds, circumventing the inherent trade-off. When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments. In case we detect the OOD setting, we employ the invariant model. We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model. The results can be found in Table 2. By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data. Evidently, the novel environment detection only works with set summaries. A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 2.
4.5 Experiment 4: Violated Criteria
Setup
We consider the PACS dataset [34], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing. The dataset includes images with labels that we intend to predict. Regarding a second classification task, we delve into the OfficeHome dataset [55]. In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
(a) Environment Art in PACS dataset. The environment is almost completely inferable from one input sample (Criterion 2.2 not satisfied). Conclusively our approach does not yield benefits.
(b) Environment Product in OfficeHome dataset. Although the environment is not inferable from one input sample (Criterion 2.2), the environment information does not yield benefits (Criterion 2.3).
Figure 6: Experiment 4. Examples where at least one of the necessary criteria is not satisfied and our approach cannot possibly yield benefits. For experimental details, see Appendix F.
Results
When the criteria are not met, no benefits can be achieved, even in the ID setting. This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 6). We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 5(b), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$. Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix F). Since Criterion 2.2 is not fulfilled, we anticipate Criterion 2.3 also to be wrong. This is indeed the case as 5(b) depicts. Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 5(b).
On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is. Results are depicted in 5(b). We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3. However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied. As expected and depicted in 5(b), our method does not yield benefits compared to the baseline model.
4.6 Experiment 5: Failure Case Detection
Setup
Besides unfulfilled criteria, another reason why our approach might fail to reap benefits or degrade in performance is when the distribution shift requires extrapolation. This might be unattainable by the model. We demonstrate, using the BikeSharing dataset [17], that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary. For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data. Details about the dataset and pre-processing steps can be found in Appendix H. We explore four scenarios, each entailing training on all seasons except one. We aim to assess the abilities of our model compared to a baseline model in detecting novel environments.
Results
In Table 3 we demonstrate that our approach is slightly superior compared to the baseline model in the ID settings. However, both the baseline and our approach experience performance degradation in the novel environments. To detect the novel environment (and consequentially potential failure cases), we compute the score as suggested in Section 2.7 and evaluate how well it distinguishes between ID versus OOD environment. We designate an independent ID test set and use the environment excluded during training (e.g., summer or winter) as the OOD set for evaluation. The area under the ROC-curve (AUROC) in Table 3 demonstrates that the score based on the summary vector provided by the permutation invariant network allows for a perfect novel environment detection whereas the standard approach fails in detecting the novel environment.
Table 3: Experiment 5. Performance comparison between our model and the baseline, broken down by target domain. We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC). Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case. We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
5 Conclusions
In this work, we introduce a novel approach enabling the extraction and utilization of contextual information from set inputs within a Domain Generalization (DG) setting. The set inputs originate from the same environment as the sample for which a prediction is required and are summarized into a vector via a permutation-invariant network. We formulate criteria that are necessary for our approach to yield benefits and are easy to verify. Empirically, we demonstrate that we can verify these criteria, which enables us to identify cases where our approach does not yield benefits. We showcase the merits of our approach on several datasets. Additionally, we demonstrate that novel environments can be detected, allowing for the identification of potential failure cases.
In our current framing, we have not yet employed regularization techniques, such as enforcing the set vector to contain all information about the environment in our composite model in Figure 1. In general, investigating inductive biases for the composite network architecture is an interesting future research avenue. Finally, even though we focused our experiments in standard supervised learning settings, our method could also be employed in other realms, such as domain disentanglement, data generation, or in combination with other DG methods. Thus, extending our method to different settings represents another exciting research direction.
Acknowledgments
JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation. We thank Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
Appendix A Theory
A.1 Insufficiency of Criteria 2 and 3 for Criteria 1
Criterion 2.2 and Criterion 2.3 are not sufficient to imply Criterion 2.1. This can be seen in an example with three environments $j\in\{1,2,3\}$. Assume the first two have completely identical input distributions. We presume that both input distributions adhere to a uniform distribution $\mathcal{U}[a,b]$. Furthermore, we assume that the third input distribution also follows a uniform distribution that is slightly shifted, i.e. $\mathcal{U}[a+\frac{a+b}{2},b+\frac{a+b}{2}]$. Due to the overlap between the third and the first two environments, a set input provides additional information about $E$ compared to a single sample $X$, verifying Criterion 2.2.
Regarding the mechanism relating inputs to outputs, we assume that on $[a,\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ differs, e.g., two constant functions with distinct values. We further assume that on $(\frac{a+b}{2},b+\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ does not vary with the environment, e.g., is constant. This aligns with Criterion 2.3: if we know the environment, we can improve the prediction, specifically on $[a,\frac{a+b}{2}]$.
However, Criterion 2.1 is not satisfiable. The set input allows us to distinguish environment 3 (i.e. the one with support $\mathcal{U}[a+\frac{a+b}{2},b+\frac{a+b}{2}]$) from the other ones. Yet, we cannot distinguish between environment 1 and environment 2. Since the relation between $X$ and output $Y$ differs only in the supports of environment 1 and environment 2 (specifically, it differs in $\mathcal{U}[a,\frac{a+b}{2}]$), the set input cannot provide additional information about the output $Y$ compared to the single input $X$, i.e. it holds $Y\perp\mathbf{S}^{(n)}\mid X$,
A.2 Proof of Theorem
In the following, we give proofs of Theorem 2.1 (a) - (d).
Proof.
For the upcoming proofs, we extensively employ the chain rule of mutual information:
$I(\mathbf{Y};Z,\mathbf{X})=I(\mathbf{Y};Z\mid\mathbf{X})+I(\mathbf{Y};\mathbf{X})$ (8)
Additionally, we have the inequalities $I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})\leq I(\mathbf{Y};E\mid\mathbf{X})$ and $I(\mathbf{S}^{(n)};\mathbf{Y}\mid\mathbf{X})\leq I(E;\mathbf{Y}\mid\mathbf{X})$ that follow from the data processing inequality and how $\mathbf{S}^{(n)}$ relates to the other variables (see Figure 1).
For (b): We easily achieve
$I(Y;\mathbf{S}^{(n)},\mathbf{X})$ $=I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})+I(\mathbf{Y};\mathbf{X})$ (9)
$\leq I(\mathbf{Y};E\mid\mathbf{X})+I(\mathbf{Y};\mathbf{X})$ (10)
$=I(\mathbf{Y};\mathbf{X})$ (11)
Therefore, we have
$0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{Y};\mathbf{S}^{(n)},\mathbf{X})-I(\mathbf{X};\mathbf{Y})\leq 0$ (12)
which proves (a).
For (a): We can write
$I(\mathbf{S}^{(n)};\mathbf{Y},\mathbf{X})$ $=I(\mathbf{S}^{(n)};\mathbf{Y}\mid\mathbf{X})+I(\mathbf{S}^{(n)};\mathbf{X})$ (13)
$\leq I(\mathbf{S}^{(n)};E\mid\mathbf{X})+I(\mathbf{S}^{(n)};\mathbf{X})$ (14)
$=I(\mathbf{S}^{(n)};\mathbf{X})$ (15)
and therefore
$0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{S}^{(n)};\mathbf{Y},\mathbf{X})-I(\mathbf{X},\mathbf{S}^{(n)})\leq 0$ (16)
and conclusively $\mathbf{Y}\perp\mathbf{S}^{(n)}\mid\mathbf{X}$.
For (c) is easily seen that $0<I(\mathbf{Y};E\mid\mathbf{X})=I(\mathbf{Y};g(\mathbf{S}^{(n)})\mid\mathbf{X})\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})$ and therefore (c) holds true.
For (d), we also employ the entropy $h(\mathbf{X})$ as well as the conditional entropy $h(\mathbf{X}\mid\mathbf{Y})$. We first establish that $I(A+B;C)\leq I(A;C)+I(B;C)$ for any RVs $A,B,C$ with $A\perp B$ and $A\perp B\mid C$:
$I(A+B;C)$ $=h(A+B)-h(A+B\mid C)$
=$\overset{(\star)}{=}\left(h(A)+h(B)-h(A\mid A+B)\right)-\left(h(A\mid C)+h(B\mid C)-h(A\mid A+B,C)\right)$
$=I(A;C)+I(B;C)-h(A\mid A+B)+h(A\mid A+B,C)$
$\overset{(\star\star)}{\leq}I(A;C)+I(B;C)$ (17)
()($\star$)( ⋆ ) follows with the chain rule for entropy
$h(A,A+B)$ $=h(A)+h(A+B\mid A)$ (18)
$=h(A)+h(B\mid A)\overset{A\perp B}{=}h(A)+h(B)$ (19)
$=h(A+B)+h(A\mid A+B)$ (20)
which implies $h(A+B)=h(A)+h(B)-h(A\mid A+B)$ and equally when conditioning on $C$.
[…]
Section A.2 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
$I(A+B;C\mid D)\leq I(A;C\mid D)+I(B;C\mid D)$ (21)
Since $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X}$ and $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$, we achieve
$0<I(\mathbf{Y};E\mid\mathbf{X})$ $=I(Y;g(\mathbf{S}^{(n)})+Z\mid\mathbf{X})$ (22)
$\leq I(\mathbf{Y};g(\mathbf{S}^{(n)})\mid\mathbf{X})+I(\mathbf{Y};Z\mid\mathbf{X})$ (23)
$\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})+I(\mathbf{Y};Z\mid\mathbf{X})$ (24)
and therefore
$0<I(\mathbf{Y};E\mid\mathbf{X})-I(\mathbf{Y};Z\mid\mathbf{X})\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})$ (25)
which concludes the proof. ∎
In the following, we discuss the assumptions in (c) and (d). In our experiments, we observed that in most datasets a relatively small sample size suffices to infer the environment label with approximately 100% accuracy (see Table 5). Therefore, the assumption that there exists a function $g(\mathbf{S}^{(n)})=E$ seems justified if $n$ is sufficiently large. To generalize the assumption where the environment label is not fully inferable, we have to make assumptions. For one, we require $\mathbf{S}^{(n)}\mid Z\mid\mathbf{X}$. This can be interpreted as “increasing the set size does not improve the prediction of $E$”. Also $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$ can be interpreted similarly: increasing the set size and considering the ground truth label/value does not enhance the predictability of $E$. Both assumptions should hold approximately if $n$ is large enough. With the assumption $I(\mathbf{Y};E\mid\mathbf{X})>I(Z;\perp\mathbf{Y}\mid\mathbf{X})$ we assume that the noise $Z$ is less predictive of $Y$ compared to $E$ if $\mathbf{X}$ is given. This can be roughly interpreted as the noise does not prove useful for predicting $Y$ from $\mathbf{X}$ compared to the ground truth environment label.
Appendix B Experiments: General Remarks
We define the relative improvements $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ as
$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})}$ (26)
and
$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})}$ (27)
$\mathcal{R}_{\text{II}}$ signifies the relative performance gain in predicting the environment when the set input is given compared to the solitude input. In contrast, $\mathcal{R}_{\text{III}}$ denotes the relative performance improvement of the environment-oracle model compared to the baseline model.
Due to the large amount of settings, we did only little hyper-parameter optimization (we looked into batch size, learning rate, and network size). For a given dataset we optimized only on one scenario where an environment is left out during training. The found hyper-parameters were then applied on all other scenarios. To ensure that the baseline model is comparable to ours, we ensure that the inference network (and feature extractor) in Figure 1 have a comparable number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$. For an overview, see Table 5. Throughout all experiments, we employ a mean-pooling operation.
We show the accuracies of classifying the environment of the contextual-environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ and the baseline environment model $f^{E\mid\mathbf{X}}$ in Table 5. Here we only consider the datasets where we performed a full evaluation of all criteria.
Appendix C Experiment 1: Details
C.1 Data Generation
Simpson’s Paradox [45, 57] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined. There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [6] or an evaluation of the efficacy of different treatments for kidney stones [9].
In order to replicate this, we create a dataset as a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target. Unless otherwise specified, we generate the data by taking an equal number of samples from each mixture component, defining the environment as a one-hot vector over the mixture components.
The mixture components are chosen to lie on a trend line that is opposite to the trend within each mixture. We achieve this by using a negative global trend, and choosing the covariance matrix of each mixture as a scaled and rotated identity matrix with opposite trend.
Setting Value Controls
n_domains $5$ number of mixture components
n_samples $10000$ number of samples per mixture component
spacing $2.0$ spacing between means of the mixture components
noise $0.25$ overall noise level
noise_ratio $6.0$ ratio of the primary to secondary noise axis
rotation_range $(45.0,45.0)$ min (leftmost) and max (rightmost) mixture rotation angle
Table 4: Default Settings for the Simpson’s Paradox Dataset. Samples from the dataset constructed with these settings can be seen in Figure 2
The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [38].
C.2 Training Details
We consider five distinct settings, where in each setting, one domain is left out during training, and considered for evaluation as a novel environment. To gauge the uncertainty stemming from data sampling, we also consider five dataset seeds for partitioning into training, validation, and test sets. For each dataset seed and model, we consider the results due to the best performance on the validation set.
We enforced that our approach and the baseline model have a similar amount of parameters for the feature extractor and final inference model. We conducted minimal hyperparameter tuning (focusing on parameters such as the learning rate schedule, batch size, and the number of parameters), and this was performed solely within one “leave-one-environment-out” setting. In total, we trained the five models outlined in Table 1 using five distinct dataset seeds. Consequently, a total of $5\cdot 5\cdot 5=125$ models were trained. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$. We choose the mean as the pooling operation.
C.3 Non-Linear Models
In the experiments in Section 4.2, we considered linear models for our model and the baseline. In the following, we show results for the non-linear model class in Figure 7. We compare predictions of a baseline model and our model on all environments in Figure 8. We see that the extrapolation task fails in some cases as in environment 1. This is due to the mismatch of the considered model class and ground truth model.
Figure 7: Experiment 1. Verification of criteria. In I we depict the relative improvement of our approach versus a baseline model. We also show I (OOD) on OOD data. In II we show the relative improvement of the oracle model compared to the baseline. In III we compare the relative improvement of the contextual environment model with respect to the baseline environment model.
Figure 8: Experiment 1. Models are trained on all environments except the OOD environment. “Extrapolation”, i.e. when environment 1 or 5 is OOD, is a particularly hard task in this setting. The set-based model shows slightly better extrapolation capabilities. Generally, our model exhibits adaptability to diverse environments, addressing a limitation present in the baseline model.
Appendix D Experiment 2: Details
(a) Environment 1
(b) Environment 2
(c) Environment 3
(d) Environment 4
Figure 9: Experiment 2. We generate four distinct domains synthetically. Notably, the background color within each domain follows a normal distribution. However, there are variations in the means across these domains Note that there is a huge overlap between the environments.
Data samples from different environments are depicted in Figure 9. The process of how inputs relate to outputs is described in Section 4.3.
During training, we employ a convolutional network to extract features $g(\mathbf{X})$. These features are passed to the inference network and the set-encoder. The feature extractor is then jointly trained with the inference network and set-encoder. We ensured that the feature extractor plus inference network and the baseline model have a comparable amount of parameters. The set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$. As a pooling operation we choose the mean-pooling.
Appendix E Experiment 3: Details
To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data. Therefore, we follow the approach proposed in Section 2.7. We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$. Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension. To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see Section 2.7 for details).
In total, we explore five dataset seeds to partition into training, validation, and test sets. To train an invariant model, we considered the same split in training, validation, and test set where the background color has no association with the label. Therefore the invariant model learns to ignore the background color and only utilize the shape for prediction. To learn effectively about the environment, we considered a large set input, namely 1024 samples in $\mathbf{S}^{(n)}$. We employed a simple set-encoder incorporating a mean pooling operation.
Appendix F Experiment 4 and 5: Details
For the BikeSharing dataset we consider a simple feed-forward neural network in all models. For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized. Here, we employ the Clip features proposed in [47]. The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions. Given that Clip features considerably simplify the task, we performed a minimal hyper-parameter search and ensured that the inference model had a similar number of parameters as the baseline model. In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
Dataset / Set Size Simpson / $32$
Domain 1 2 3 4 5
$f^{E\mid\mathbf{X}}$ $86.3\pm 1.3$ $90.8\pm 1.3$ $90.7\pm 0.8$ $90.4\pm 0.9$ $85.5\pm 0.8$
$f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$
Dataset / Set Size ProDAS / $128$ OfficeHome / $4$ PACS / $4$
Domain 1 2 3 4 Product Art
$f^{E\mid\mathbf{X}}$ $43.8\pm 1.1$ $50.0\pm 1.3$ $49.9\pm 2.3$ $44.4\pm 1.0$ $86.16\pm 0.33$ $\mathbf{99.72}\pm 0.33$
$f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ $\mathbf{99.6}\pm 0.6$ $\mathbf{99.5}\pm 1.0$ $\mathbf{98.7}\pm 1.6$ $\mathbf{98.0}\pm 3.2$ $\mathbf{98.49}\pm 0.24$ $\mathbf{100.0}\pm 0.0$
Table 5: Environment classification accuracy for different models and datasets, broken down by domain. As in Table 3, the uncertainty (mean and standard deviation) is computed over multiple seeds for dataset splits. In all cases, the set-based model outperforms the baseline.
In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
Appendix G Comparison of Permutation-Invariant Architectures
As a pilot experiment, we estimate the contextual information contained in a set input by evaluating the binary classification accuracy of a set-based model compared to a baseline model with singleton sample input.
Importantly, we postulate that for stronger domain overlap, the contextual information contained within the single sample decreases significantly, while the contextual information within the set decreases only weakly, depending on the set size. Domains that do not overlap exactly will remain distinguishable, so long as the set size is large enough.
Therefore, we construct the toy dataset as described in Section C.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
We train each architecture on this dataset for $20$ epochs, using $5$ different seeds. We evaluate a total of $30$ domain spacings, linearly distributed between $0.05$ and $1.5$ (both inclusive). Since we evaluate a baseline model, plus 3 set-based models at 3 different set sizes, this brings us to a total of $30\cdot 20\cdot 5\cdot(1+3\cdot 3)=30000$ model epochs. We choose the batch size at $128$ fixed.
Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [1] activation. For the set-based methods, this is followed by the respective pooling. We choose $8$ heads for the attention-based model.
Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [73].
For methods that support a non-unit output set size, we choose the output set size as $4$. The output set is mean-pooled prior to projection into the logit space.
Figure 10: Comparison of different architectural choices for the permutation-invariant network in predicting the data’s originating environment. We consider various distances between environments and different set sizes $n$. As anticipated, the plots illustrate that smaller environment distances make it more challenging to differentiate between them. Moreover, with larger set size $n$, our ability to predict the environment label improves. Notably, the baseline model shows significantly poorer performance compared to the model utilizing contextual information in the form of a set input.
Appendix H Bike Sharing Dataset
This dataset, taken from the UCI machine learning repository[17], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.
Each dataset entry contains information about the season, time, and weather at the time of rental. Casual renters are also distinguished from registered ones.
Similar to [49], we only consider the hourly rental data. We drop information about the concrete date and information about casual versus registered renters. We choose the season variable (spring, summer, fall, winter) as the environment and the bike rental count as the regression target. Since we deal with count data, we also apply square root transformation on the target similar to [49].

### v2

@@
-Towards Context-Aware Domain Generalization: Representing Environments with Permutation-Invariant Networks
+Towards Context-Aware Domain Generalization: Understanding the Benefits and Limits of Marginal Transfer Learning
 Abstract
-In this work, we show that information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models when applied in new domains or production environments.
-We formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same environment/domain as the input itself.
-These representations are jointly learned with a standard supervised learning objective, providing incremental information about the unknown outcome.
-Furthermore, we offer a theoretical analysis of the conditions under which our approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice.
-Additionally, we contribute insights into the kind of distribution shifts for which our approach promises robustness.
-Our empirical evaluation demonstrates the effectiveness of our approach for both low-dimensional and high-dimensional data sets.
+In this work, we analyze the conditions under which information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models in new domains.
+Following work in marginal transfer learning in Domain Generalization (DG), we formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same domain as the input itself.
+We offer a theoretical analysis of the conditions under which this approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice.
+Additionally, we contribute insights into the kind of distribution shifts for which the marginal transfer learning approach promises robustness.
+Empirical analysis shows that our criteria are effective in discerning both favorable and unfavorable scenarios.
 Finally, we demonstrate that we can reliably detect scenarios where a model is tasked with unwarranted extrapolation in out-of-distribution (OOD) domains, identifying potential failure cases.
@@
 We assume that the environment $E$ is a source node that is not caused by any system variable and that the relationship between $\mathbf{X}$ and $\mathbf{Y}$ varies with the environment.
-$\mathbf{S}^{(n)}$ is a set of $n$ IID inputs available in the new environment (i.e.
+$\mathbf{S}^{(n)}$ is a set of $n$ i.i.d.
+inputs available in the new environment (i.e.
 context).
-B) The source component shift corresponding to the assumed DGP and example data coinciding with Simpson’s paradox.
-C) The workings of our approach in a test environment.
+B) The workings of the context-aware model (aka the marginal transfer learning approach) in a test environment.
 A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context.
@@
 The set-representation can be combined with the input to reliably detect out-of-distribution (OOD) queries and prevent failure cases in domain generalization due to model misspecification.
-Distribution shifts are the cause of many failure cases in machine learning [27, 32] and the root of various peculiar phenomena in classical statistics, such as Simpsons’ paradox [45, 57].
-In this work, we employ permutation-invariant neural networks as set-encoders [15, 7] to improve the predictions of standard supervised models under distribution shift.
-Specifically, we consider training in the realm of Domain Generalization (DG), a setting where data from distinct environments [^1] We use the terms environment and domain interchangeably.
-is available for training and testing [39, 74].
-For illustration, consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$.
+Distribution shifts are the cause of many failure cases in machine learning [Hendrycks and Dietterich, 2019, Koh et al., 2021] and the root of various peculiar phenomena in classical statistics, such as Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021].
+The Domain Generalization (DG) task seeks a model that is robust to distribution shifts by utilizing data from distinct environments during training [Muandet et al., 2013, Zhou et al., 2022].
+In the context of DG, marginal transfer learning enhances a model with context information to achieve better predictions Blanchard et al.
+[2021].
+The “context” of a test instance is simply a set of samples that stems from the same environment as the instance itself and can be embedded, for instance, by permutation-invariant neural networks Edwards and Storkey [2017], Bloem-Reddy and Teh [2020].
+In this work, we enhance the fundamental understanding of settings where marginal transfer learning in DG can reap benefits compared to baseline models.
+Consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$.
 Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands.
 It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$.
-This raises two questions:
-  • (1)
-    Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$?
-  • (2)
-    How should $E$ be represented to maximize the performance gain?
-The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$.
+This raises two questions: (i) Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$ and (ii) how should $E$ be represented to maximize the performance gain?
+The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$ alone.
 For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region.
 Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically.
-A straightforward answer to the second question is to distinguish environments by discrete labels.
-However, we argue that $E$ should be a continuous embedding.
-First, continuous embeddings can also be computed for new environments that have not been present in the training data.
-Second, when $p(\mathbf{Y}\mid\mathbf{X},E)$ receives a continuous $E$, it can potentially configure itself for unseen environments by interpolating between the training environments.
-And finally, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
-In the current work, we systematically investigate both questions, formalize three criteria when $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and demonstrate how continuous embeddings $E$ can be learned from auxiliary data by means of set encoders (see Figure 1).
+A straightforward answer to the second question is to distinguish environments by discrete labels, but we propose to learn continuous embeddings instead.
+This allows us to compute representations for novel environments that enable teal-time adaptation and interpolation between the training environments.
+Moreover, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
+Building on previous work in marginal transfer learning [Blanchard et al., 2021], we learn a continuous embedding of $E$ from auxiliary data using set encoders, as depicted in Figure 1.
+To systematically address the first question, we establish three criteria that delineate the circumstances in which $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and subsequently prove their necessity.
 Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach.
-First, we require that a single input alone is insufficient to deduce the originating environment (see Criterion 2.2).
-If this condition is not met, our model cannot possibly extract more information about the originating environment, compared to a standard model.
-Second, we require that, if the environment label is known by an oracle and given as an additional input to the standard model, its performance must improve (see Criterion 2.3).
-Otherwise, the set-based representation of the environment cannot possibly yield benefits.
-Notably, these two criteria are easy to verify and also met under the source component shift (see Section 2.6), which is a dataset shift that occurs in many scenarios [46].
-When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and the potential to generate silent failures.
-While our approach is not exempt from this “curse of extrapolation”, it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [41].
-Moreover, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs.
+When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and an increased risk of silent failures.
+While marginal transfer learning is not exempt from this “curse of extrapolation”, we find that it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [Müller et al., 2023].
+Accordingly, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs.
 models that are robust to out-of-distribution (OOD) scenarios on the fly.
-Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [65, 42, 37].
-Accordingly, we can adaptively select the most robust model in the OOD setting and the most predictive model in the ID setting, an approach we demonstrate on the ColoredMNIST data set [4].
+Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [Yang et al., 2022b, Müller et al., 2022, Magliacane et al., 2018].
 In summary, our contributions are:
-  • We propose a novel approach to Domain Generalization (DG) that leverages context information from new environments in the form of learnable set-representations;
-  • We formalize the necessary and empirically verifiable conditions under which our approach can reap benefits from context information and improve on standard approaches;
-  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required.
+  • We formalize the necessary and empirically verifiable conditions under which the marginal transfer learning approach can reap benefits from context information and improve on standard approaches;
+  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required;
+  • We show how the detection of novel environments allows for model selection, overcoming the trade-off between predictive performance and robustness.
 2 Method
+In the following, we will briefly introduce context-aware models (aka marginal transfer learning) in DG.
+We then discuss criteria under which we can expect to reap benefits from the additional set-representation.
+Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which context-aware approaches may prove advantageous in practice.
+Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
 2.1 Notation
 We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively.
-We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ an i.i.d.
-sample (i.e., a set of further inputs) from the given domain.
+We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ a set of $n$ i.i.d.
+samples (i.e., a set of further inputs) from the given domain.
 The domain label $E$ is only known during training time and unknown during inference.
-2.2 General Idea
-The key idea of our approach is to build models that utilize not only a singleton input $\mathbf{X}$ to predict a target $\mathbf{Y}$, but also information about the environment of $\mathbf{X}$ that can improve the prediction $\hat{\mathbf{Y}}$.
-Providing environmental information in the form of a one-hot label is hardly feasible, as it presupposes the exact number of possible environments to be known during training, and that we always know from which environment an input originates at inference time.
-Consequently, such an approach is doomed to fail when the test data originates from a novel or unknown environment.
-To overcome this problem, we employ permutation-invariant neural networks to adaptively represent environmental information given a set of test inputs.
-We will first detail our approach and then discuss criteria under which we can expect to reap benefits from the additional set-representation.
-Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which our approach may prove advantageous in practice.
-Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
-2.3 Permutation-Invariant Neural Networks
-As mentioned above, a basic goal of our approach is to synthesize contextual information about a target input $\mathbf{X}$ by compressing a set of $n$ further inputs $\mathbf{S}^{(n)}:=\{\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n}\}$ from the same environment into a permutation-invariant representation.
-The notion of permutation invariance is closely related to a core concept in probabilistic modeling and Bayesian inference – exchangeability [44].
-Accordingly, an exchangeable sequence of random vectors is characterized by a joint distribution which is invariant to any permutation of the elements:
-$p(\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n})=p(\mathbf{X}_{\pi(1)},\mathbf{X}_{\pi(2)},\dots,\mathbf{X}_{\pi(n)}),$ (1)
-where $\pi:\mathbb{N}\rightarrow\mathbb{N}$ denotes an arbitrary permutation of index elements $n\in\mathbb{N}$.
-Exchangeable observations may come in various forms, for instance, patients entering a hospital, different measurements obtained with the same device, or sets of visual images, such as faces in a crowd.
-When it comes to learning exchangeable symmetries, permutation-invariant functions can serve as a key building block in neural architectures [7], as they automatically encode a favorable inductive bias towards permutation-invariance by design.
-A simple and intuitive way to create permutation-invariant functions involves the sum-decomposition
-$h(\mathbf{S}^{(n)})=\rho\left(\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})\right),$ (2)
-where $\sigma$ and $\rho$ can be any functions, including deep neural networks [69, 7].
-$h$ is permutation-invariant due to the summation operator which ensures that the argument of $\rho$ is agnostic to the order of elements in the set $\mathbf{S}^{(n)}$.
-Despite having favorable theoretical properties, plain sum-decompositions can have limited representational capacity in practice [69, 58].
-Thus, more expressive permutation-invariant functions can be learned by stacking equivariant transformations with sum-decompositions [69, 7] or by using a self-attention mechanism without positional encodings [33].
-In order to re-assess the expressiveness of these methods, we show a comparison of the binary domain classification accuracy based on domain overlap and set size in Figure 10.
-2.4 Context-Aware Model
-Our model consists of two key components (also illustrated in Figure 1):
-  • a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and
-  • an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction.
+2.2 Context-Aware Models
+A context-aware model consists of two key components (also illustrated in Figure 1): (i) a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and (ii) an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction.
 The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short.
 For a given supervised learning task, we aim to find the minimum to the following optimization problem
-$\boldsymbol{\theta}^{\star}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{p(\mathbf{X},\mathbf{Y},E)}\left[c(f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)}),\mathbf{Y})\right],$ (3)
-where $c$ is a task-specific loss function (e.g., cross-entropy for classification or mean squared error for regression).
-The algorithmic details for optimizing Equation 3 are detailed in Algorithm 1.
-For practical reasons, we first apply a feature extractor $g$ and then pass $\{g(\mathbf{X})\mid\mathbf{X}\in\mathbf{S}^{(n)}\}$ as input to the set encoder and $g(\mathbf{X})$ as input to the inference network, building upon the features extracted by $g$.
-Note, that $g$ can be a pre-trained network, in which case we can treat it as a fixed transformation, or its parameters can be optimized jointly with $\boldsymbol{\theta}$.
-Data: Samples from the joint distribution $p(\mathbf{X},\mathbf{Y},E)$
-Input : Composite model parameters $\boldsymbol{\theta}$, set size $n$, batch size $m$, loss-function $c$, number of iterations $k$, learning rate schedule $\alpha(k)$
-1 for $i=1,\dots,k$ do
-2       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
-3       for $j=1,\dots,m$ do
-4             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
-5             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
-6       end for
-7      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
-$\boldsymbol{\theta}_{k}\leftarrow\boldsymbol{\theta}_{k-1}-\alpha(k)\nabla_{\boldsymbol{\theta}}\left(\sum_{j=1}^{m}c\left(f_{\boldsymbol{\theta}}(\mathbf{x}_{j},\mathbf{s}_{j}^{(n)}),\boldsymbol{y}_{j}\right)\right)$
-8 end for
-Output : Trained context-aware model $f_{\boldsymbol{\theta}}$
-Algorithm 1 Optimizing Equation 3 for context-aware domain generalization.
+$\widehat{\boldsymbol{\theta}}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{p(\mathbf{X},\mathbf{Y},E)}\left[c(f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)}),\mathbf{Y})\right],$ (1)
+where $c$ is a task-specific loss function (e.g., cross-entropy for classification or mean squared error for regression). Algorithm 1 details the optimization ofEquation 1.
 2.3 Criteria for Improvement
-In the following, we establish criteria under which our method can exploit the distribution shifts between environments and yield improved predictions.
+In the following, we establish criteria under which context information allows to exploit the distribution shifts between environments and yield improved predictions.
 In total, we propose three criteria that are necessary to achieve incremental improvement.
@@
 Criterion 2.1.
-$\mathbf{S}^{(n)}\not\perp\mathbf{Y}\mid\mathbf{X}$ or equivalently $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})>0$.
+$\mathbf{S}^{(n)}\not\perp\mathbf{Y}\mid\mathbf{X}$ or $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})>0$.
 The second criterion requires that, given a target input $\mathbf{X}$, a set of further i.i.d.
@@
 Criterion 2.2.
-$E\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ or equivalently $I(E;\,\mathbf{S}^{(n)}\mid\mathbf{X})>0$.
+$E\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ or $I(E;\,\mathbf{S}^{(n)}\mid\mathbf{X})>0$.
 In Figure 1, an instance $\mathbf{X}$ cannot be assigned with complete certainty to an environment.
@@
 Crucially, this criterion is not satisfied, if we can recover the origin environment from the singleton input $\mathbf{X}$ alone.
-The third criterion requires that the singleton input $\mathbf{X}$ gains information about $\mathbf{Y}$ if we also consider the origin environment $E$ of $\mathbf{X}$.
+The third criterion requires that the singleton input $\mathbf{X}$ carries information about $\mathbf{Y}$ if we also consider the origin environment $E$ of $\mathbf{X}$.
 Criterion 2.3.
-$\mathbf{Y}\not\perp E\mid\mathbf{X}$ or equivalently $I(\mathbf{Y};\,E\mid\mathbf{X})>0$.
-In Figure 1, this is evidently the case: If we knew the environment from which the data stems, we could improve our prediction of $\mathbf{Y}$.
-Furthermore, this criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy.
+$\mathbf{Y}\not\perp E\mid\mathbf{X}$ or $I(\mathbf{Y};\,E\mid\mathbf{X})>0$.
+This criterion can serve as a sanity check in case we have an oracle that can identify the origin environment of the data with perfect accuracy.
 In what follows, we show that Criterion 2.2 and Criterion 2.3 are necessary conditions for Criterion 2.1.
 We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1.
-We even generalize this result for the case where the environment label is not inferable with 100% accuracy.
+We generalize this result for the case where the environment label is not inferable with 100% accuracy in Appendix C.
 Theorem 2.1.
@@
 This conveys that if we could perfectly infer $E$ from $\mathbf{S}^{(n)}$, then Criterion 2.3 implies Criterion 2.1.
-  • (d)
-    Assume that there exists a function $g$ and a noise variable $Z$ that elicits the relation $E=g(\mathbf{S}^{(n)})+Z$ and satisfies $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X}$ as well as $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$.
-Furthermore, assume that $\mathbf{Y}\not\perp E\mid\mathbf{X}$ and $I(\mathbf{Y};E\mid\mathbf{X})>I(Z;\mathbf{Y}\mid\mathbf{X})$.
-Then, we achieve $\mathbf{Y}\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$, recovering Criterion 2.1.
-The proof for this theorem and a discussion of its presumptions can be found in Section A.2.
+The proof for this theorem can be found in subsection C.3.
 Unfortunately, we cannot conclude that $Y\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general.
-An example where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in Section A.1.
+A counterexample where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in subsection C.2.
+It is also worth noting that Criterion 2.3 might be achievable while Criterion 2.2 is unattainable and vice versa.
+For instance, when we can infer the originating environment from one sample (Criterion 2.2 is not attainable), the relation between $\mathbf{X}$ and $\mathbf{Y}$ might still vary with the environment (Criterion 2.3 is achievable).
 2.4 Source Component Shift
 Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satifsied.
-Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [46, Chapter 1.9].
+Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [Quinonero-Candela et al., 2008].
 The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$.
@@
 First, the input distribution changes whenever the environment changes.
-[^2] In our case, we require a stronger assumption, namely, that the domains are not deducible from a single input, as formalized in Criterion 2.2.
+[^1] In our case, we need a stronger assumption: the domains cannot be deduced from a single input, as defined in Criterion 2.2.
 Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1).
-For more details on this kind of distribution shift, we refer the reader to [46, Chapter 1.9].
-It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [45, 57], which supplies a proof-of-concept for our approach (see Experiment 1).
+For more details on this kind of distribution shift, we refer the reader to Quinonero-Candela et al.
+[2008].
+It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021], which supplies a proof-of-concept for our approach (see Experiment 1).
+An important point to highlight is that the frequently encountered covariate shift where only $P(\mathbf{X})$ in $P(\mathbf{X},\mathbf{Y})=P(\mathbf{Y}\mid\mathbf{X})P(\mathbf{X})$ varies between environments [Quinonero-Candela et al., 2008], does not conform to the conditions specified in Criterion 2.3.
+Hence, context-aware models do not provide advantages when compared to standard models under covariate shift (under the assumption all models converged to the optimum).
 2.5 Detection of Novel Environments
@@
 In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation.
-Following [41], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$.
+Following Müller et al.
+[2023], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$.
 As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder.
 Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment.
-Following the approach in [41], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment.
+Following the approach in Müller et al.
+[2023], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment.
 To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set.
-We also compare our novel environment detector with the same score function computed solely from singleton features $g(\mathbf{X})$.
-These results can be quickly previewed in Table 2.
+We also compare our novel environment detector with the same score function computed from singleton features $g(\mathbf{X})$ alone (see Table 1 for a preview).
 3 Related Work
 3.1 Domain Generalization
-The aim of domain generalization (DG) is to train models that generalize well under distribution shifts [39, 74].
-The DG setting involves access to data from multiple domains during training exploiting the heterogeneity between domains.
-In this context, the plethora of algorithms aimed at improving robustness has been divided into three categories [60]: (i) algorithms that enhance the data (e.g., via data augmentation or domain synthesis [56, 68]), (ii) algorithms that adapt the learning algorithm (e.g., self-supervised learning [29, 8]), and (iii) algorithms aiming to find robust representations (e.g., via invariance learning [4, 42]).
-Our approach explores a different avenue exploiting contextual information about the origin of the data via an adaptive environment embedding.
-In contrast to domain adaptation (DA) [61], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training.
-As a middle-ground between DA and DG, test-time adaptation (TTA) involves the provision of unlabeled samples during test time, enabling further fine-tuning of the model [35].
-TTA can be categorized into test-time domain adaptation, test-time batch adaptation (TTBA), and online test-time adaptation (OTTA) [35].
-Among these settings, our work aligns most closely with the TTBA scenario concerning the utilization of environment information.
-In TTBA a pre-trained model is adapted to one or a few inputs [54, 72, 51].
-Each adaptation depends on the mini-batch at hand.
-Similarly, we consider a mini-batch as a set input to deliver contextual information.
-However, we do not adapt or fine-tune our model at test time, but rather directly extract context information via the set-encoder.
-This allows us to identify whether extrapolation is necessary, enabling model misspecification detection [50].
-Moreover, since we do not need to fine-tune the model at test time, our approach is considerably more efficient.
-Finally, [12] assume a setting where inputs from multiple domains are available, but it is unknown which sample belongs to which environment – even during training time.
-The authors infer potential domain labels that are used for downstream invariant learning, for example, via Invariant Risk Minimization (IRM) [4].
-This method improves on baseline models and enables the application of DG methods in the absence of domain labels.
-Interestingly, it could be combined with our approach when no environment labels are provided during training, and we leave this as an avenue for future research.
+The aim of Domain Generalization (DG) is to train models that generalize well under distribution shifts [Muandet et al., 2013, Zhou et al., 2022].
+In contrast to domain adaptation (DA) [Wang and Deng, 2018], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training.
+Both, DG and DA, involve access to data from multiple domains during training exploiting the heterogeneity between domains.
+For an overview on DG see for instance Wang et al.
+[2022], Zhou et al.
+[2022].
+As a middle-ground between DA and DG, test-time adaptation (TTA) [Liang et al., 2023] and marginal transfer learning [Blanchard et al., 2021] involves the provision of unlabeled samples during test time.
+In TTA, the model is typically fine-tuned according to the provided samples [Liang et al., 2023].
+Marginal transfer learning in DG refers to the setting where access to the marginal feature distribution is given during testing and provided to the model via $\frac{1}{n}\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})$ [Blanchard et al., 2021].
+In this context, $\sigma$ represents a feature extraction model applied to each $\mathbf{X}_{i}$, and the summation ensures that the operation remains invariant under permutations.
+Different choices for $\sigma$ are explored in the literature.
+Specifically, Zhang et al.
+[2021] utilizes CNNs, Blanchard et al.
+[2021], Dubey et al.
+[2021] incorporates kernel embeddings, and Bao and Karaletsos [2023] adopts patch embeddings.
+While these works highlight the advantages of their methodologies, they fall short in analyzing the exact settings and conditions that contribute to these benefits compared to a standard model.
+Additionally, they do not address the detection of potential failure cases or the utilization of context embedding for model selection.
+A recent method proposes to use transformers to exploit previously seen samples instead of learning permutation-invariant embedding [Gupta et al., 2023]
+The trade-off between ID performance and OOD performance is well-known [Müller et al., 2022, Magliacane et al., 2018, Zhang et al., 2023].
+A method to mitigate this trade-off was proposed in Zhang et al.
+[2023].
+Their approach presupposes knowledge of the data’s originating environment, whereas our goal is to infer it.
 3.2 Learning Permutation-Invariant Representations
-Analyzing set-structured data with neural networks has received much theoretical [58, 7, 43] and empirical [69, 33, 70] momentum in recent years.
-For instance, [70] build on the set transformer architecture [33] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling.
-The resulting “PICASO” blocks include consecutive multi-head attention stacks with skip connections, which update their template(s) depending on previous template(s) and the particular input set (in contrast to global templates).
-Differently, [13] propose to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
-The authors also show how to use their criterion for generative and few-shot classification tasks In a somewhat similar vain, [5] investigate a gradient-based optimization method for aggregating set-structured data.
-Importantly, these methods comprise a pool of algorithms which can be used as the backbone architecture for realizing the set-encoder in our approach.
+Analyzing set-structured data with neural networks has received much theoretical [Wagstaff et al., 2022, Bloem-Reddy and Teh, 2020, Murphy et al., 2018] and empirical [Zaheer et al., 2017, Lee et al., 2019, Zare and Van Nguyen, 2021] momentum in recent years.
+For instance, Zare and Van Nguyen [2021] build on the set transformer architecture [Lee et al., 2019] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling.
+Differently, dan Guo et al.
+[2021] proposes to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
+Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks.
+In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [Orbanz and Roy, 2014].
+Typically, hierarchical (aka multi-level) Bayesian models are used to deal with Simpson’s paradox by organizing observations into clusters or levels [Wikle, 2003, Gelman et al., 2013], mirroring the notions of domain or environment in DG.
+Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
+The machine learning literature is rich with variational methods for learning invariant summary statistics from set-structured data [Edwards and Storkey, 2016, Garnelo et al., 2018, Kim et al., 2019, 2021, Zeng et al., 2022].
+Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations.
+To achieve this goal, we employ variants of the DeepSet [Zaheer et al., 2017] and SetTransformer [Lee et al., 2019] architectures for our backbone set-encoder throughout all experiments.
+Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
+Figure 2: Experiment 1.
+Left: Toy dataset that conforms to our theoretical criteria.
+A detailed description can be found in subsection E.1.
+Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3.
+Marginal distributions of the five environments are shown on top.
+Right: Comparison of environment classification accuracy for different set sizes.
+Distances between environments refer to the distance between the means of the environments.
+3.3 OOD Detection and Selective Classification
+Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [Aggarwal and Yu, 2001, Yang et al., 2021, Shen et al., 2021, Han et al., 2022, Yang et al., 2022a].
+Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [Yang et al., 2021].
+OOD detection is closely related to inference with a reject option (also termed selective classification) [Geifman and El-Yaniv, 2017, El-Yaniv et al., 2010], which allows classifiers to refrain from predicting ambiguous or novel conditions [Hendrickx et al., 2021].
+The reject option has been extensively studied in statistical and machine learning [Hellman, 1970, Fumera and Roli, 2002, Grandvalet et al., 2008, Wegkamp and Yuan, 2012], with early work dating back to the 1950s [Chow, 1957, 1970, Hellman, 1970].
+More recently, Müller et al.
+[2023] explored the utility of selective classification in DG settings.
+They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent.
+In this work, we consider a post-hoc score based on the $k$-nearest neighbours to the training set in feature space similar to Sun et al.
+[2022], which applies to both classification and regression settings.
+Unlike the approach taken in Müller et al.
+[2023], where the focus lies on features of individual instances, we consider the set summary provided by the set-encoder.
+Thus, we can identify novel environments even when singleton inputs lack sufficient information.
+4 Experiments
+In the following, we explore various aspects of the context-aware model across three different dimensions.
+First, we show on two datasets that a context-aware model achieves improved performance in the ID as well as the OOD setting compared to a baseline model when the conditions of a source component shift are met.
+Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model.
+We also show that novel environment detection can be utilized to avoid failure cases.
+Third, we demonstrate that the necessary criteria (see subsection 2.3) can be validated empirically, identifying cases where no benefits of the method can be expected.
+Experimental details can be found in the Appendix.
+Furthermore, in Appendix B, we conduct an additional experiment resembling Experiment 1, but with high-dimensional inputs
+Figure 3: Experiment 1.
+Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on a toy example.
+We also show I (OOD) on OOD data.
+II depicts the relative improvement of the environment-oracle model compared to the baseline model.
+III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it.
+Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
+4.1 Evaluation Approach
+To approximate Criterion 2.1, Criterion 2.2 and Criterion 2.3, we are required to train five distinct models (see Table 3 for an overview).
+We denote our composite model as $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ (see Figure 1) and the baseline model (having no access to the context) as $f^{\mathbf{Y}\mid\mathbf{X}}$.
+Based on these two models, we can compute the relative improvement achieved by our model relative to a baseline model via
+$\mathcal{R}_{\text{I}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (2)
+Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$, both measured on an unseen test set.
+$\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by the context-aware approach and therefore the fulfillment of Criterion 2.1.
+In the regression setting, we consider the negative L2-Loss as the performance measure.
+To validate Criterion 2.2, we train a contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$.
+Additionally, we train a baseline environment model $f^{E\mid\mathbf{X}}$ aimed at predicting $E$ solely from $\mathbf{X}$.
+We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^2] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix D of the contextual environment predictor relative to the baseline environment predictor.
+$\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied.
+In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
+Similarly, we consider an environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ that aims to predict $\mathbf{Y}$ from the singleton input and the environment label $E$.
+We define the relative improvement $\mathcal{R}_{\text{III}}$ of the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ compared to the baseline method $f^{\mathbf{Y}\mid\mathbf{X}}$.
+In this case, the relative improvement $\mathcal{R}_{\text{III}}$ is associated with Criterion 2.3.
+4.2 Experiment 1: Toy Example
+Setup
+To set the stage, we consider a dataset shown in Figure 2 and explained in detail in subsection E.1.
+The dataset includes data from five different environments, defined by distinct Gaussian distributions.
+Each Gaussian deviates due to its location (i.e.
+mean vector).
+The dataset exemplifies Simpson’s paradox, wherein a naive fit of the data (without accounting for environmental factors) would yield a negatively sloped line (see Figure 10).
+Conversely, using environmental information for prediction enables to capture the true underlying model with positively sloped lines, counteracting the overall negative trend of the mean vectors.
+Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X$, obtained by projecting the entire sample onto the $x$-axis in Figure 2.
+Thus, the setting aligns with Criterion 2.2, and corresponds to Criterion 2.3, since knowing the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
+Results
+As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input.
+Figure 2 illustrates that additional set input improves the ability to distinguish between environments significantly and the more samples we include, the better the distinction.
+As expected, a decrease in the distance between environment marginal means necessitates more samples to differentiate between environments.
+Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix I.
+Figure 4: Experiment 1. Predictions performed on the toy dataset illustrated in Figure 2. We show predictions made by both our set-encoder approach and the vanilla model in the ID and OOD settings.
+Next, we assess the predictive capabilities of the context-aware approach across all possible scenarios of “leave-one-environment-out”.
+This involves training on all environments except one and treating the excluded environment as a novel OOD scenario.
+Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in subsection E.3).
+We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3.
+Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3).
+We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I).
+This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself.
+Note that for III we achieve less relative improvement since we consider the accuracy here and not the L2-Loss.
+Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4 for one trained model.
+Our model captures and utilizes the characteristics of each environment for prediction.
+In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences.
+Note that we obtained the best results by considering a class of linear models that aligns with the data-generating process.
+However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see subsection E.3).
+4.3 Experiment 2: Colored MNIST
+Setup
+The ColoredMNIST dataset [Arjovsky et al., 2019] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1).
+Furthermore, label noise is deliberately added, so only in 75% of all cases, the label can correctly be predicted from the shape.
+To make things more challenging, the image background can take two colors that are also associated with the image label.
+In the first environment, the association is 90% and in the second one 80%.
+Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape.
+However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
+This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [Arjovsky et al., 2019, Zhang et al., 2023].
+For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%).
+In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
+[…]
+ID OOD
+Baseline $\mathbf{84.6}\pm 0.3$ $10.2\pm 0.3$
+Invariant $72.8\pm 0.9$ $\mathbf{73.1}\pm 0.2$
+Selection (Ours) $\mathbf{84.1}\pm 0.3$ $\mathbf{73.1}\pm 0.2$
+Selection (Baseline) $\mathbf{84.0}\pm 0.3$ $14.0\pm 0.4$
+Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
+Table 1: Experiment 2.
+Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs.
+The features extracted by our model allow for improved OOD detection compared to the features of the baseline model.
+Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
+Results
+Here, we assume the invariant model to be given (see Appendix G for details), but it could also be obtained by invariant learning, e.g.
+Invariant Risk Minimization [Arjovsky et al., 2019].
+With our novel environment detection approach (see subsection 2.5) we can get the best of both worlds, circumventing the inherent trade-off.
+When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments.
+In case we detect the OOD setting, we employ the invariant model.
+We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model.
+The results can be found in Table 1.
+By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data.
+Evidently, the novel environment detection only works with set summaries.
+A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 1.
+4.4 Experiment 3: Violated Criteria
+Setup
+We consider the PACS dataset [Li et al., 2017], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing.
+The dataset includes images with labels that we intend to predict.
+Regarding a second classification task, we delve into the OfficeHome dataset [Venkateswara et al., 2017].
+In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
+(a) Environment Art in PACS dataset.
+The environment is almost completely inferable from one input sample (Criterion 2.2 not satisfied).
+Conclusively the context-aware approach does not yield benefits.
+(b) Environment Product in OfficeHome dataset. Although the environment is not inferable from one input sample (Criterion 2.2), the environment information does not yield benefits (Criterion 2.3).
+Figure 5: Experiment 3.
+Tell-tale examples where at least one of the necessary criteria is not satisfied and the context-aware approach cannot possibly yield benefits.
+For experimental details, see Appendix H.
+Results
+When the criteria are not met, no benefits can be achieved, even in the ID setting.
+This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 5).
+We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 4(a), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$.
+Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix H).
+Also Criterion 2.3 is not satisfied as 4(a) depicts.
+Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 4(a).
+On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is.
+Results are depicted in 4(b).
+We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3.
+However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied.
+As expected and depicted in 4(b), our method does not yield benefits compared to the baseline model.
+4.5 Experiment 4: Failure Case Detection
+Setup
+Besides unfulfilled criteria, another reason why a context-aware approach might fail to reap benefits is when the distribution shift requires extrapolation.
+This might be unattainable by the model, making the inclusion of a reject option beneficial.
+Using the BikeSharing dataset [Fanaee-T, 2013], we demonstrate that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary.
+For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data.
+Here we explore the scenario where we train on all seasons except winter.
+Details about the dataset, pre-processing steps and other scenarios can be found in Appendix J.
+Results
+In Table 5 we demonstrate that the context-aware approach is slightly superior compared to the baseline model in the ID settings.
+However, both the baseline and the context-aware approach experience performance degradation in the novel winter environment.
+To detect the novel environment and, consequentially, potential failure cases, we compute the score as suggested in subsection 2.5 and evaluate how well it distinguishes between ID versus OOD environments.
+We designate an independent ID test set and use the environment excluded during training (here winter) as the OOD set for evaluation.
+The area under the ROC-curve (AUROC) in Table 5 demonstrates that the score based on the permutation-invariant embedding allows for a perfect detection of the novel environment, whereas the standard approach fails as expected.
+Table 2: Experiment 4.
+Performance comparison between our model and the baseline for the winter domain in the BikeSharing dataset.
+We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC).
+Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case.
+We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
+Other domains can be found in Appendix J.
+5 Conclusions
+In this work, we aimed to advance the theoretical understanding of marginal transfer learning in domain generalization.
+Thus, we formulated criteria that are necessary for the approach to yield benefits and are also easy to verify empirically.
+Moreover, we pinpoint the source component shift as a scenario where context-aware models can offer advantages, enabling the identification of favorable scenarios.
+Additionally, we illustrate the capability to identify novel environments, thereby allowing for the detection of potential failure cases.
+This, in turn, also supports model selection by choosing the most predictive on ID data and the most robust model on OOD data.
+Acknowledgments
+JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation. We thank Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
+Appendix A Algorithm
+Data: Samples from the joint distribution $p(\mathbf{X},\mathbf{Y},E)$
+Input : Composite model parameters $\boldsymbol{\theta}$, set size $n$, batch size $m$, loss-function $c$, number of iterations $k$, learning rate schedule $\alpha(k)$
+for $i=1,\dots,k$ do
+       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
+       for $j=1,\dots,m$ do
+             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
+             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
+       end for
+      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
+$\boldsymbol{\theta}_{k}\leftarrow\boldsymbol{\theta}_{k-1}-\alpha(k)\nabla_{\boldsymbol{\theta}}\left(\sum_{j=1}^{m}c\left(f_{\boldsymbol{\theta}}(\mathbf{x}_{j},\mathbf{s}_{j}^{(n)}),\boldsymbol{y}_{j}\right)\right)$
+end for
+Output : Trained context-aware model $f_{\boldsymbol{\theta}}$
+Algorithm 1 Optimizing Equation 1 for context-aware domain generalization.
+A.1 Model Overview
+For a detailed overview of the models used to verify the criteria, see Table 3.
 Model Symbol Description Purpose
@@
 Table 3: The five different model types used to evaluate our approach and verify the theoretical criteria for improvement.
-Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks.
-In contrast, probabilistic models attempt to learn a conditional or a marginal distribution over the set-representations.
-In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [44].
-In particular, hierarchical or multi-level Bayesian models are used to model the dependencies in nested data, where observations are organized into clusters or levels [64, 21], mirroring the notions of domain or environment in DG.
-Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
-From a somewhat different Bayesian perspective, IID data represents the starting point for learning invariant summary statistics for parameter estimation or model comparison.
-The pioneering work on Neural Statisticians [14] tackles the task from a variational perspective, learning global set-representations as part of a generative model.
-Neural processes [19, 30] comprise a related family of set-based models for prediction and uncertainty quantification in supervised learning tasks, rooted in the spirit of Gaussian processes [48].
-Finally, [31] explore a permutation-invariant variational autoencoder (SetVAE) with multiple latent variables trained with a modified ELBO objective.
-The goal of this and follow-up models [71] is accurate generative performance, less so the learning of compact representations.
-The theoretical and empirical implications of learning random set-representations falls outside the scope of our current work, but offers a potentially fruitful avenue for future research.
-Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations.
-For the purpose of achieving this goal, we employ variants of the DeepSet [69] and SetTransformer architectures for our backbone set-encoder throughout all experiments.
-Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
-3.3 OOD Detection and Selective Classification
-Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [3, 67, 52, 23, 66].
-Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [67].
-OOD detection is closely related to inference with a reject option (also termed selective classification) [20, 16], which allows classifiers to refrain from making a prediction under ambiguous or novel conditions [25].
-The reject option has been extensively studied in statistical and machine learning [24, 18, 22, 62], with early work dating back to the 1950s[11, 10, 24].
-More recently, [41] explored the utility of selective classification in DG settings.
-They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent.
-Post-hoc methods rely on various aspects of pre-trained model outputs, such as the softmax outputs (e.g., [28]), logit outputs (e.g., [26, 36]), or intermediate features (e.g., [2, 53, 59]).
-In this work, we consider a post-hoc score based on the $k$-nearest neighbors to the training set in feature space similar to [53], which is applicable to both classification and regression settings.
-Unlike the approach taken in [41], where the focus lies on individual input features, we consider the set summary provided by the set-encoder.
-Thus, we can identify novel environments even when singleton inputs lack sufficient information.
-4 Experiments
-Figure 2: Experiment 1.
-Left: Toy dataset that conforms to our theoretical criteria.
-A detailed description can be found in Section C.1.
-Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3.
-Marginal distributions of the five environments are shown on top.
-Right: Comparison of environment classification accuracy for different set sizes.
-In the following, we explore various aspects of our approach across three different dimensions.
-First, we show on two datasets that our model achieves improved performance in the ID as well as the OOD setting compared to a baseline model.
-Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model.
-We also show that novel environment detection can be utilized to avoid failure cases.
-Third, we demonstrate that the necessary criteria (see Section 2.5) can be validated empirically, identifying cases where no benefits of our method can be expected.
-Experimental details can be found in the Appendix.
-4.1 Evaluation Approach
-To approximate Criterion 2.1, Criterion 2.2 and Criterion 2.3, we are required to train five distinct models (see Table 1 for an overview).
-We denote our composite model as $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ (see Figure 1) and the baseline model (having no access to the context) as $f^{\mathbf{Y}\mid\mathbf{X}}$.
-Based on these two models, we can compute the relative improvement achieved by our model relative to a baseline model via
-$\mathcal{R}_{\text{I}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (4)
-Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$.
-$\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by our approach and therefore the fulfillment of Criterion 2.1.
-In the regression setting, we consider the negative L2-Loss as the performance measure.
-To validate Criterion 2.2, we train a contextual environment model (referred to as $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$) utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$.
-Additionally, we train a baseline environment model (denoted as $f^{E\mid\mathbf{X}}$) aimed at predicting $E$ solely from $\mathbf{X}$.
-We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^3] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix B of the contextual environment predictor relative to the baseline environment predictor.
-$\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied.
-In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
-Similarly, we consider an environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ that aims to predict $\mathbf{Y}$ from the singleton input and the environment label $E$.
-We define the relative improvement $\mathcal{R}_{\text{III}}$ of the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ compared to the baseline method $f^{\mathbf{Y}\mid\mathbf{X}}$.
-In this case, the relative improvement $\mathcal{R}_{\text{III}}$ is associated with Criterion 2.3.
-4.2 Experiment 1: Toy Example
+Appendix B Additional Experiment: ProDAS
 Setup
-To set the stage, we consider a dataset that is inspired by [63].
-The dataset includes data from five environments, defined by five 2D Gaussian distributions differing only in their locations (i.e., mean vectors).
-The data-generating process is thus:
-$j$ $\sim\text{Categorical}(p_{1},\dots,p_{J})$ (5)
-$\mathbf{X}^{(ij)}$ $\sim\mathcal{N}(\boldsymbol{\mu}^{(j)},\mathbf{\Sigma})\quad\text{for}\quad i=1,\dots,I_{j}$ (6)
-This is the classical setting of Simpsons’ paradox, for which a naive linear model fit oblivious to the hierarchical data-generating process will yield results opposite to the true relationship between $X_{1}^{(j)}$ and $X_{2}^{(j)}$ in each environment (cf.
-Figure 2 and Section C.1 for details).
-We designate the first dimension, $X_{1}$, as the input feature, while the second dimension, $Y=X_{2}$, serves as the outcome variable.
-Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X_{1}$, obtained by projecting the entire sample onto the $x$-axis in Figure 2.
-Thus, this setting aligns with Criterion 2.2, and, additionally, corresponds to Criterion 2.3, since learning about the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
-Figure 3: Experiment 1.
-Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on toy example.
-We also show I (OOD) on OOD data.
-II depicts the relative improvement of the environment-oracle model compared to the baseline model.
-III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it.
-Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
-Results
-As a first check of Criterion 2.2, we evaluate whether a set input provides additional information about the environment compared to a singleton input.
-This evaluation involves studying the classification accuracy in distinguishing between the environments.
-In Figure 2 we observe that the additional set input improves the ability to predict the environment significantly and the more samples we include, the better the prediction.
-As expected, a decrease in the distance between environments necessitates more samples to differentiate between environments.
-Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix G.
-Next, we assess the predictive capabilities of our approach across all possible scenarios of “leave-one-environment-out”.
-This involves training on all environments except one and treating the excluded environment as a novel OOD scenario.
-Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in Section C.3).
-We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3.
-Providing contextual information in the form of a set input increases the performance significantly compared to a baseline model in the ID as well as in the OOD setting (see I and I (OOD) in Figure 3).
-We also observe a slightly higher relative improvement when the environment label is directly provided (see II) compared to using the output of the set-encoder (see I).
-This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself.
-Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4.
-Our model captures and utilizes the characteristics of each environment for prediction.
-In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences.
-Note that we obtained the best results on this problem by considering a class of linear models that aligns with the data-generating process.
-However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see Section C.3).
-Figure 4: Experiment 1. Predictions performed on the toy dataset illustrated in Figure 2. We show predictions made by both our set-encoder approach and the vanilla model in the ID and OOD settings.
-4.3 Experiment 2: ProDAS Example
-Setup
-We utilize the ProDAS library [40] to generate high-dimensional image data that meets our dataset requirements.
+We utilize the ProDAS library [Müller et al., 2023] to generate high-dimensional image data that meets our dataset requirements.
 The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size.
 Additionally, the background varies in color and texture, resulting in a complex scenario.
+For examples see Figure 9.
 We consider the task of predicting the object size.
@@
 Notably, inferring the originating environment from a single sample is unattainable due to overlapping background distributions (corresponding to Criterion 2.2).
-Samples of different environments are shown in Appendix D.
+Samples of different environments are shown in Appendix F.
 This example could be inspired by microscopy data where different microscopes correspond to distinct environments, each exhibiting its own characteristics.
@@
 All formal criteria are satisfied and a very significant improvement is achieved, both in the ID and the OOD setting, by considering the contextual information from the environment.
-Details for this experiment can be found in Appendix D.
-4.4 Experiment 3: Colored MNIST
-Setup
-The ColoredMNIST dataset [4] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1).
-Furthermore, there exists label noise, so only in 75% of all cases, the label can correctly be predicted from the input image.
-To make things more challenging, the image background can take two colors that are also associated with the image label.
-In the first environment, the association is 90% and in the second one 80%.
-Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape.
-However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
-This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [42].
-For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%).
-In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
+Additional details for this experiment can be found in Appendix F.
 Figure 6: Experiment 2: Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means, no improvement is achieved) on ProDAS dataset.
@@
 Variations arise from using different seeds to partition the ID data into training, test and validation set.
-[…]
-In-Distribution Out-of-Distribution
-Baseline $\mathbf{84.6}\pm 0.3$ $10.2\pm 0.3$
-Invariant $72.8\pm 0.9$ $\mathbf{73.1}\pm 0.2$
-Selection (Ours) $\mathbf{84.1}\pm 0.3$ $\mathbf{73.1}\pm 0.2$
-Selection (Baseline) $\mathbf{84.0}\pm 0.3$ $14.0\pm 0.4$
-Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
-Table 2: Experiment 3.
-Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs.
-The features extracted by our model allow for improved OOD detection compared to the features of the baseline model.
-Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
-Results
-Here, we assume the invariant model to be given (see Appendix E for details), but it could also be obtained by invariant learning, e.g.
-Invariant Risk Minimization [4].
-With our novel environment detection approach (see Section 2.7) we can get the best of both worlds, circumventing the inherent trade-off.
-When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments.
-In case we detect the OOD setting, we employ the invariant model.
-We compare this kind of model selection due to the features $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ inherent to our model versus the features extracted by the baseline model.
-The results can be found in Table 2.
-By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data.
-Evidently, the novel environment detection only works with set summaries.
-A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 2.
-4.5 Experiment 4: Violated Criteria
-Setup
-We consider the PACS dataset [34], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing.
-The dataset includes images with labels that we intend to predict.
-Regarding a second classification task, we delve into the OfficeHome dataset [55].
-In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
-(a) Environment Art in PACS dataset. The environment is almost completely inferable from one input sample (Criterion 2.2 not satisfied). Conclusively our approach does not yield benefits.
-(b) Environment Product in OfficeHome dataset. Although the environment is not inferable from one input sample (Criterion 2.2), the environment information does not yield benefits (Criterion 2.3).
-Figure 6: Experiment 4. Examples where at least one of the necessary criteria is not satisfied and our approach cannot possibly yield benefits. For experimental details, see Appendix F.
-Results
-When the criteria are not met, no benefits can be achieved, even in the ID setting.
-This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 6).
-We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 5(b), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$.
-Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix F).
-Since Criterion 2.2 is not fulfilled, we anticipate Criterion 2.3 also to be wrong.
-This is indeed the case as 5(b) depicts.
-Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 5(b).
-On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is.
-Results are depicted in 5(b).
-We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3.
-However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied.
-As expected and depicted in 5(b), our method does not yield benefits compared to the baseline model.
-4.6 Experiment 5: Failure Case Detection
-Setup
-Besides unfulfilled criteria, another reason why our approach might fail to reap benefits or degrade in performance is when the distribution shift requires extrapolation.
-This might be unattainable by the model.
-We demonstrate, using the BikeSharing dataset [17], that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary.
-For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data.
-Details about the dataset and pre-processing steps can be found in Appendix H.
-We explore four scenarios, each entailing training on all seasons except one.
-We aim to assess the abilities of our model compared to a baseline model in detecting novel environments.
-Results
-In Table 3 we demonstrate that our approach is slightly superior compared to the baseline model in the ID settings.
-However, both the baseline and our approach experience performance degradation in the novel environments.
-To detect the novel environment (and consequentially potential failure cases), we compute the score as suggested in Section 2.7 and evaluate how well it distinguishes between ID versus OOD environment.
-We designate an independent ID test set and use the environment excluded during training (e.g., summer or winter) as the OOD set for evaluation.
-The area under the ROC-curve (AUROC) in Table 3 demonstrates that the score based on the summary vector provided by the permutation invariant network allows for a perfect novel environment detection whereas the standard approach fails in detecting the novel environment.
-Table 3: Experiment 5.
-Performance comparison between our model and the baseline, broken down by target domain.
-We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC).
-Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case.
-We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
-5 Conclusions
-In this work, we introduce a novel approach enabling the extraction and utilization of contextual information from set inputs within a Domain Generalization (DG) setting.
-The set inputs originate from the same environment as the sample for which a prediction is required and are summarized into a vector via a permutation-invariant network.
-We formulate criteria that are necessary for our approach to yield benefits and are easy to verify.
-Empirically, we demonstrate that we can verify these criteria, which enables us to identify cases where our approach does not yield benefits.
-We showcase the merits of our approach on several datasets.
-Additionally, we demonstrate that novel environments can be detected, allowing for the identification of potential failure cases.
-In our current framing, we have not yet employed regularization techniques, such as enforcing the set vector to contain all information about the environment in our composite model in Figure 1.
-In general, investigating inductive biases for the composite network architecture is an interesting future research avenue.
-Finally, even though we focused our experiments in standard supervised learning settings, our method could also be employed in other realms, such as domain disentanglement, data generation, or in combination with other DG methods.
-Thus, extending our method to different settings represents another exciting research direction.
-Acknowledgments
-JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation. We thank Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
-Appendix A Theory
-A.1 Insufficiency of Criteria 2 and 3 for Criteria 1
+Appendix C Theory
+C.1 Generalization of Theorem 2.1 to Noisy Environments
+Theorem C.1.
+In addition to Theorem 2.1, the following holds:
+  • (d)
+    Assume that there exists a function $g$ and a noise variable $Z$ that elicits the relation $E=g(\mathbf{S}^{(n)})+Z$ and satisfies $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X}$ as well as $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$.
+Furthermore, assume that $\mathbf{Y}\not\perp E\mid\mathbf{X}$ and $I(\mathbf{Y};E\mid\mathbf{X})>I(Z;\mathbf{Y}\mid\mathbf{X})$.
+Then, we achieve $\mathbf{Y}\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$, recovering Criterion 2.1.
+The proof can be found in subsection C.3.
+C.2 Insufficiency of Criteria 2 and 3 for Criterion 1
 Criterion 2.2 and Criterion 2.3 are not sufficient to imply Criterion 2.1.
@@
 it holds $Y\perp\mathbf{S}^{(n)}\mid X$,
-A.2 Proof of Theorem
+C.3 Proof of Theorem 2.1
 In the following, we give proofs of Theorem 2.1 (a) - (d).
@@
 $0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{Y};\mathbf{S}^{(n)},\mathbf{X})-I(\mathbf{X};\mathbf{Y})\leq 0$ (8)
-which proves (a).
+which proves (b).
 For (a): We can write
@@
 and therefore
-$0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{S}^{(n)};\mathbf{Y},\mathbf{X})-I(\mathbf{X},\mathbf{S}^{(n)})\leq 0$ (16)
+$0\leq I(\mathbf{Y};\mathbf{S}^{(n)}\mid\mathbf{X})=I(\mathbf{S}^{(n)};\mathbf{Y},\mathbf{X})-I(\mathbf{X};\mathbf{S}^{(n)})\leq 0$ (12)
 and conclusively $\mathbf{Y}\perp\mathbf{S}^{(n)}\mid\mathbf{X}$.
@@
 […]
-Section A.2 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
+subsection C.3 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
 $I(A+B;C\mid D)\leq I(A;C\mid D)+I(B;C\mid D)$ (17)
@@
 This can be roughly interpreted as the noise does not prove useful for predicting $Y$ from $\mathbf{X}$ compared to the ground truth environment label.
-Appendix B Experiments: General Remarks
+Appendix D Experiments: General Remarks
 We define the relative improvements $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ as
-$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})}$ (26)
+$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X}})}$ (22)
 and
-$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})}$ (27)
+$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}$ (23)
 $\mathcal{R}_{\text{II}}$ signifies the relative performance gain in predicting the environment when the set input is given compared to the solitude input.
@@
 For a given dataset we optimized only on one scenario where an environment is left out during training.
-The found hyper-parameters were then applied on all other scenarios.
+The found hyper-parameters were then applied to all other scenarios.
 To ensure that the baseline model is comparable to ours, we ensure that the inference network (and feature extractor) in Figure 1 have a comparable number of parameters as the baseline model.
@@
 Here we only consider the datasets where we performed a full evaluation of all criteria.
-Appendix C Experiment 1: Details
-C.1 Data Generation
-Simpson’s Paradox [45, 57] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined.
-There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [6] or an evaluation of the efficacy of different treatments for kidney stones [9].
-In order to replicate this, we create a dataset as a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target.
+Appendix E Experiment 1: Details
+E.1 Data Generation
+Simpson’s Paradox [Peters et al., 2017, von Kügelgen et al., 2021] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined.
+There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [Bickel et al., 1975] or an evaluation of the efficacy of different treatments for kidney stones [Charig et al., 1986].
+To replicate this, we create a dataset inspired by an illustration of Simpson’s Paradox on Wikipedia sim .
+The dataset consists of a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target.
 Unless otherwise specified, we generate the data by taking an equal number of samples from each mixture component, defining the environment as a one-hot vector over the mixture components.
 The mixture components are chosen to lie on a trend line that is opposite to the trend within each mixture.
-We achieve this by using a negative global trend, and choosing the covariance matrix of each mixture as a scaled and rotated identity matrix with opposite trend.
+We achieve this by using a negative global trend and choosing the covariance matrix of each mixture as a scaled and rotated identity matrix with an opposite trend.
 Setting Value Controls
@@
 Table 4: Default Settings for the Simpson’s Paradox Dataset. Samples from the dataset constructed with these settings can be seen in Figure 2
-The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [38].
-C.2 Training Details
+The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [minutephysics and Reich, 2017].
+E.2 Training Details
 We consider five distinct settings, where in each setting, one domain is left out during training, and considered for evaluation as a novel environment.
@@
 We choose the mean as the pooling operation.
-C.3 Non-Linear Models
-In the experiments in Section 4.2, we considered linear models for our model and the baseline.
+E.3 Non-Linear Models
+In the experiments in subsection 4.2, we considered linear models for our model and the baseline.
 In the following, we show results for the non-linear model class in Figure 7.
@@
 Generally, our model exhibits adaptability to diverse environments, addressing a limitation present in the baseline model.
-Appendix D Experiment 2: Details
+Appendix F Additional Experiment: Details
 (a) Environment 1
@@
 (d) Environment 4
-Figure 9: Experiment 2.
+Figure 9: Additional Experiment.
 We generate four distinct domains synthetically.
@@
 However, there are variations in the means across these domains Note that there is a huge overlap between the environments.
-Data samples from different environments are depicted in Figure 9. The process of how inputs relate to outputs is described in Section 4.3.
+Data samples from different environments are depicted in Figure 9. The process of how inputs relate to outputs is described in Appendix B.
 During training, we employ a convolutional network to extract features $g(\mathbf{X})$.
@@
 As a pooling operation we choose the mean-pooling.
-Appendix E Experiment 3: Details
+Appendix G Experiment 2: Details
 To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data.
-Therefore, we follow the approach proposed in Section 2.7.
+Therefore, we follow the approach proposed in subsection 2.5.
 We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$.
 Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension.
-To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see Section 2.7 for details).
+To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see subsection 2.5 for details).
 In total, we explore five dataset seeds to partition into training, validation, and test sets.
@@
 We employed a simple set-encoder incorporating a mean pooling operation.
-Appendix F Experiment 4 and 5: Details
+Appendix H Experiment 3 and 4: Details
 For the BikeSharing dataset we consider a simple feed-forward neural network in all models.
 For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized.
-Here, we employ the Clip features proposed in [47].
+Here, we employ the Clip features proposed in Radford et al.
+[2021].
 The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions.
@@
 In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
+Table 5: Experiment 4.
+Performance comparison between our model and the baseline, broken down by target domain.
+We compare their performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC).
+Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case.
+We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
 Dataset / Set Size Simpson / $32$
@@
 In all cases, the set-encoder is kept simple and its hyper-parameters are selected for optimal performance of the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$.
-Appendix G Comparison of Permutation-Invariant Architectures
+Appendix I Comparison of Permutation-Invariant Architectures
 As a pilot experiment, we estimate the contextual information contained in a set input by evaluating the binary classification accuracy of a set-based model compared to a baseline model with singleton sample input.
@@
 Domains that do not overlap exactly will remain distinguishable, so long as the set size is large enough.
-Therefore, we construct the toy dataset as described in Section C.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
+Therefore, we construct the toy dataset as described in subsection E.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
 We train each architecture on this dataset for $20$ epochs, using $5$ different seeds.
@@
 We choose the batch size at $128$ fixed.
-Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [1] activation.
+Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [Agarap, 2019] activation.
 For the set-based methods, this is followed by the respective pooling.
 We choose $8$ heads for the attention-based model.
-Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [73].
+Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [Zhang and Sabuncu, 2018].
 For methods that support a non-unit output set size, we choose the output set size as $4$. The output set is mean-pooled prior to projection into the logit space.
@@
 As anticipated, the plots illustrate that smaller environment distances make it more challenging to differentiate between them.
-Moreover, with larger set size $n$, our ability to predict the environment label improves.
+Moreover, with a larger set size $n$, our ability to predict the environment label improves.
 Notably, the baseline model shows significantly poorer performance compared to the model utilizing contextual information in the form of a set input.
-Appendix H Bike Sharing Dataset
-This dataset, taken from the UCI machine learning repository[17], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.
+Appendix J Bike Sharing Dataset
+This dataset, taken from the UCI machine learning repository [Fanaee-T, 2013], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.
 Each dataset entry contains information about the season, time, and weather at the time of rental. Casual renters are also distinguished from registered ones.
-Similar to [49], we only consider the hourly rental data.
+Similar to Rothenhäusler et al.
+[2021], we only consider the hourly rental data.
 We drop information about the concrete date and information about casual versus registered renters.
 We choose the season variable (spring, summer, fall, winter) as the environment and the bike rental count as the regression target.
-Since we deal with count data, we also apply square root transformation on the target similar to [49].
+Since we deal with count data, we also apply square root transformation on the target similar to Rothenhäusler et al.
+[2021].

### v3

@@
-Towards Context-Aware Domain Generalization: Understanding the Benefits and Limits of Marginal Transfer Learning
+Reliable ML From Unreliable Data
 Abstract
-In this work, we analyze the conditions under which information about the context of an input $\mathbf{X}$ can improve the predictions of deep learning models in new domains.
-Following work in marginal transfer learning in Domain Generalization (DG), we formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same domain as the input itself.
+In this work, we analyze the conditions under which information about the context of an input data point can improve the predictions of deep learning models in new domains.
+Following work in marginal transfer learning and domain generalization, we formalize the notion of context as a permutation-invariant representation of a set of data points that originate from the same domain as the input itself.
 We offer a theoretical analysis of the conditions under which this approach can, in principle, yield benefits, and formulate two necessary criteria that can be easily verified in practice.
@@
 1 Introduction
+Distribution shifts are the cause of many failure cases in machine learning [1, 2] and the root of various peculiar phenomena in classical statistics, such as Simpson’s paradox [3, 4].
+Domain Generalization (DG) seeks models that are robust to distribution shifts by utilizing data from distinct environments during training [5, 6].
+In the context of DG, marginal transfer learning enhances a model with context information to achieve better predictions [7].
+The “context” of a test instance is a set of samples that stems from the same environment as the instance itself and can be embedded, for instance, by permutation-invariant neural networks [8].
+In this work, we enhance the fundamental understanding of settings where marginal transfer learning in DG can reap benefits compared to baseline models.
+Consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$.
+Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands.
+It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$.
+This raises a key question: Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$?
+This question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$ alone.
+For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region.
+Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically.
 Figure 1: Conceptual sketch of our setup and approach.
@@
 $\mathbf{S}^{(n)}$ is a set of $n$ i.i.d.
-inputs available in the new environment (i.e.
-context).
-B) The workings of the context-aware model (aka the marginal transfer learning approach) in a test environment.
+inputs available in the new environment.
+The bidirectional arrow indicates that the causal relation between $\mathbf{X}$ and $\mathbf{Y}$ could be explained by a common cause or $\mathbf{Y}$ causing $\mathbf{X}$ (or vice versa).
+B) The context-aware model (marginal transfer learning approach) in a test environment.
 A set-encoder generates a permutation-invariant representation $h(\mathbf{S}^{(n)})$ of the context.
-An inference network (e.g., a classifier) processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome (e.g., label) of the target input.
-The set-representation can be combined with the input to reliably detect out-of-distribution (OOD) queries and prevent failure cases in domain generalization due to model misspecification.
-Distribution shifts are the cause of many failure cases in machine learning [Hendrycks and Dietterich, 2019, Koh et al., 2021] and the root of various peculiar phenomena in classical statistics, such as Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021].
-The Domain Generalization (DG) task seeks a model that is robust to distribution shifts by utilizing data from distinct environments during training [Muandet et al., 2013, Zhou et al., 2022].
-In the context of DG, marginal transfer learning enhances a model with context information to achieve better predictions Blanchard et al.
-[2021].
-The “context” of a test instance is simply a set of samples that stems from the same environment as the instance itself and can be embedded, for instance, by permutation-invariant neural networks Edwards and Storkey [2017], Bloem-Reddy and Teh [2020].
-In this work, we enhance the fundamental understanding of settings where marginal transfer learning in DG can reap benefits compared to baseline models.
-Consider a probabilistic model $p(\mathbf{Y}\mid\mathbf{X})$ that classifies diseases $\mathbf{Y}$ from magnetic resonance (MR) images $\mathbf{X}$.
-Since MR images are not fully standardized, the classifier should work slightly differently for images acquired by different hardware brands.
-It thus makes sense to inform the classifier about the current environment $E$ (here: hardware brand) and extend it into $p(\mathbf{Y}\mid\mathbf{X},E)$.
-This raises two questions: (i) Under which circumstances will the classifier $p(\mathbf{Y}\mid\mathbf{X},E)$ be superior to $p(\mathbf{Y}\mid\mathbf{X})$ and (ii) how should $E$ be represented to maximize the performance gain?
-The first question is important because there might exist a function $E=f(\mathbf{X})$ allowing the classifier $p(\mathbf{Y}\mid\mathbf{X})$ to deduce $E$ from the data $\mathbf{X}$ alone.
-For example, $E$ might be inferred from the periphery of the given image, while $\mathbf{Y}$ depends on its central region.
-Then, no additional information is gained by passing $E$ explicitly, and both classifiers perform identically.
-A straightforward answer to the second question is to distinguish environments by discrete labels, but we propose to learn continuous embeddings instead.
-This allows us to compute representations for novel environments that enable teal-time adaptation and interpolation between the training environments.
-Moreover, discrete and continuous $E$ are equally informative for the known environments, ensuring no loss in information.
-Building on previous work in marginal transfer learning [Blanchard et al., 2021], we learn a continuous embedding of $E$ from auxiliary data using set encoders, as depicted in Figure 1.
-To systematically address the first question, we establish three criteria that delineate the circumstances in which $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and subsequently prove their necessity.
+An inference network processes the representation along with the target input $\mathbf{X}$ and predicts the unknown outcome of the target input.
+The set-representation can be combined with the input to reliably detect out-of-distribution queries and prevent failure cases in domain generalization due to model misspecification.
+Building on previous work in marginal transfer learning [7], we aim to learn a continuous embedding of $E$ from auxiliary data using set-encoders, as depicted in Figure 1.
+We then establish three criteria that delineate the circumstances in which $p(\mathbf{Y}\mid\mathbf{X},E)$ is beneficial, and subsequently prove their necessity.
 Notably, two of these criteria are empirically testable using standard models and are shown to be necessary conditions for the success of the approach.
 When test environments are highly dissimilar to the training environments, all DG methods enter an extrapolation regime with unknown prospects of success and an increased risk of silent failures.
-While marginal transfer learning is not exempt from this “curse of extrapolation”, we find that it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [Müller et al., 2023].
-Accordingly, we propose a method to select between models that are specialized in the in-distribution (ID) setting vs.
-models that are robust to out-of-distribution (OOD) scenarios on the fly.
-Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [Yang et al., 2022b, Müller et al., 2022, Magliacane et al., 2018].
+While marginal transfer learning is not exempt from this “curse of extrapolation”, we find that it comes with a natural way to reliably detect novel environments in set-representation space and delineate its competence region [9].
+Accordingly, we propose a method to select between models that are specialized in the ID setting versus models that are robust to OOD scenarios on the fly.
+Thus, we can overcome the notorious trade-off between ID predictive performance and robustness to distribution shifts [10, 11, 12].
 In summary, our contributions are:
-  • We formalize the necessary and empirically verifiable conditions under which the marginal transfer learning approach can reap benefits from context information and improve on standard approaches;
-  • We perform an extensive empirical evaluation and show that we can reliably detect failure cases when the necessary criteria of our theory are not met, or when extrapolation is required;
+  • We formalize the necessary and empirically verifiable conditions under which the marginal transfer learning can improve on standard approaches;
+  • We show empirically that we can identify cases where context-aware models offer no advantages or when dangerous extrapolation is necessary;
   • We show how the detection of novel environments allows for model selection, overcoming the trade-off between predictive performance and robustness.
 2 Method
-In the following, we will briefly introduce context-aware models (aka marginal transfer learning) in DG.
-We then discuss criteria under which we can expect to reap benefits from the additional set-representation.
-Afterwards, we explain the theoretical data-generating process that matches these criteria, providing insight into the distribution shifts for which context-aware approaches may prove advantageous in practice.
-Finally, we discuss the process of identifying new environments that demand extrapolation, potentially leading to failure cases.
 2.1 Notation
 We denote inputs $\mathbf{X}\in\mathcal{X}$ and outputs as $\mathbf{Y}\in\mathcal{Y}$, without any strict requirements on the input and output spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively.
-We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ a set of $n$ i.i.d.
-samples (i.e., a set of further inputs) from the given domain.
-The domain label $E$ is only known during training time and unknown during inference.
+We treat the (unknown) domain label $E$ as a random variable and denote with $\mathbf{S}^{(n)}$ a set of $n$ further i.i.d.
+samples from a given domain, whose label $E$ is only known during training time.
 2.2 Context-Aware Models
-A context-aware model consists of two key components (also illustrated in Figure 1): (i) a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set encoder”) with parameters $\psi$ that maps a set input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and (ii) an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction.
+A context-aware model consists of two key components (also illustrated in Figure 1): (i) a permutation-invariant network $h_{\boldsymbol{\psi}}$ (“set-encoder”) with parameters $\psi$ that maps a set-input $\mathbf{S}^{(n)}$ to a summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, and (ii) an inference network $f_{\boldsymbol{\phi}}$ with parameters $\boldsymbol{\phi}$ that maps both the input $\mathbf{X}$ and the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to a final prediction.
 The complete model is denoted as $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})=f_{\boldsymbol{\phi}}(\mathbf{X},h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)}))$ with parameters $\boldsymbol{\theta}=(\boldsymbol{\psi},\boldsymbol{\phi})$ for short.
-For a given supervised learning task, we aim to find the minimum to the following optimization problem
+For a given supervised learning task, we consider the optimization problem
 $\widehat{\boldsymbol{\theta}}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathbb{E}_{p(\mathbf{X},\mathbf{Y},E)}\left[c(f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)}),\mathbf{Y})\right],$ (1)
@@
 The symbol $\perp$ (resp.
-⟂$\not\perp$̸) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp.
+$\not\perp$) between two random vectors $\mathbf{X}$ and $\mathbf{Y}$ is used to express that the random vectors are independent (resp.
 dependent) or conditionally independent (resp.
@@
 In general, the more data we consider, the better we can predict the originating environment.
-Crucially, this criterion is not satisfied, if we can recover the origin environment from the singleton input $\mathbf{X}$ alone.
+Crucially, this criterion is not satisfied if we can recover the origin environment from the singleton input $\mathbf{X}$ alone.
 The third criterion requires that the singleton input $\mathbf{X}$ carries information about $\mathbf{Y}$ if we also consider the origin environment $E$ of $\mathbf{X}$.
@@
 We furthermore prove that if we can extract the environment label fully from $\mathbf{S}^{(n)}$, then Criterion 2.2 and Criterion 2.3 are sufficient conditions for Criterion 2.1.
-We generalize this result for the case where the environment label is not inferable with 100% accuracy in Appendix C.
 Theorem 2.1.
@@
 This conveys that if we could perfectly infer $E$ from $\mathbf{S}^{(n)}$, then Criterion 2.3 implies Criterion 2.1.
-The proof for this theorem can be found in subsection C.3.
-Unfortunately, we cannot conclude that $Y\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general.
-A counterexample where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in subsection C.2.
-It is also worth noting that Criterion 2.3 might be achievable while Criterion 2.2 is unattainable and vice versa.
-For instance, when we can infer the originating environment from one sample (Criterion 2.2 is not attainable), the relation between $\mathbf{X}$ and $\mathbf{Y}$ might still vary with the environment (Criterion 2.3 is achievable).
+In our experiments, we observe that a function $g(\mathbf{S}^{(n)})=E$ can already be found for small $n$ (see for instance Figure 2).
+In this case, we obtain $I(\mathbf{S}^{(n)};\,\mathbf{Y}\mid\mathbf{X})=I(E;\,\mathbf{Y}\mid\mathbf{X})$ and Criterion 2.2 and Criterion 2.3 are sufficient to obtain Criterion 2.1.
+Unfortunately, we cannot conclude that $Y\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$ follows from Criterion 2.2 and Criterion 2.3 in general.
+A counterexample where Criterion 2.2 and Criterion 2.3 hold, but Criterion 2.1 is violated, is provided in Section C.2.
+We furthermore provide the proof of the theorem, an illustration for the theorem as well as a generalization of (c) in Appendix C.
+It is worth noting that model misspecification adds another layer of uncertainty when verifying the criteria.
+In cases where determining the correct mutual information is not feasible (for instance, when $p(\mathbf{Y}\mid\mathbf{X})$, $p(\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}$), or $p(\mathbf{Y}\mid\mathbf{X},E)$ cannot be learned adequately), two primary issues may emerge.
+Firstly, the effective utilization of the set-input $\mathbf{S}^{(n)}$ (or E) may be hindered due to either the restricted expressive power of the model class or a scarcity of training data.
+As a result, the context-aware model might not improve on a baseline model that only utilizes $\mathbf{X}$.
+Consequentially the criteria could seem unattainable while they actually are.
+Secondly, after training, we might observe an apparent advantage of the approximation of $p(\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)})$ or $p(\mathbf{Y}\mid\mathbf{X},E)$ over the approximation of $p(\mathbf{Y}\mid\mathbf{X})$, despite the true model not conferring any advantage.
+In this scenario, the criteria may appear to be satisfied, whereas in reality they are not.
+An example of this case can be easily constructed by considering a non-linear $p(\mathbf{Y}\mid\mathbf{X})$ and a linear function class.
 2.4 Source Component Shift
-Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satifsied.
-Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [Quinonero-Candela et al., 2008].
+Using our approach, we can characterize the kind of distribution shift that allows our criteria to be satisfied.
+Source component shift refers to the scenario where the data comes from a number of sources (or environments) each with different characteristics [13].
 The source component shift can be described by the graphical model in Figure 1, where the environment directly affects both the input $\mathbf{X}$ and the outcome $\mathbf{Y}$.
@@
 First, the input distribution changes whenever the environment changes.
-[^1] In our case, we need a stronger assumption: the domains cannot be deduced from a single input, as defined in Criterion 2.2.
 Second, the relationship between inputs and outcomes varies with the environment (corresponding to Criterion 2.1).
-For more details on this kind of distribution shift, we refer the reader to Quinonero-Candela et al.
-[2008].
-It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [Peters et al., 2017, von Kügelgen et al., 2021], which supplies a proof-of-concept for our approach (see Experiment 1).
-An important point to highlight is that the frequently encountered covariate shift where only $P(\mathbf{X})$ in $P(\mathbf{X},\mathbf{Y})=P(\mathbf{Y}\mid\mathbf{X})P(\mathbf{X})$ varies between environments [Quinonero-Candela et al., 2008], does not conform to the conditions specified in Criterion 2.3.
-Hence, context-aware models do not provide advantages when compared to standard models under covariate shift (under the assumption all models converged to the optimum).
+For more details on this kind of distribution shift, we refer the reader to [13].
+It is also worth noting that the graph in Figure 1 corresponds to Simpson’s paradox [3, 4], which supplies a proof-of-concept for our approach (see Experiment 1).
+An important point to highlight is that the frequently encountered covariate shift where only $P(\mathbf{X})$ in $P(\mathbf{X},\mathbf{Y})=P(\mathbf{Y}\mid\mathbf{X})P(\mathbf{X})$ varies between environments [13], does not conform to the conditions specified in Criterion 2.3.
+Hence, context-aware models do not provide advantages when compared to standard models under covariate shift.
 2.5 Detection of Novel Environments
@@
 In the following, we explain how we aim to detect the second case that might result in potential failure cases due to fundamental challenges in extrapolation.
-Following Müller et al.
-[2023], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$.
+Following [9], we can define a score $s(h_{\psi}(\mathbf{S}^{(n)}))$ on the summary vector $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ implicit in our model $f_{\boldsymbol{\theta}}(\mathbf{X},\mathbf{S}^{(n)})$ that aims to predict the target variable $\mathbf{Y}$.
 As a score function, we consider the distance of $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$ to the $k$-nearest neighbors in the training data in the feature space of the set-encoder.
 Accordingly, set-representations that elicit a score surpassing a certain threshold are considered to originate from a novel environment.
-Following the approach in Müller et al.
-[2023], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment.
+Following the approach in [9], we consider the score distribution and set a threshold to classify a specific percentage, denoted as $q$, of in-distribution samples as originating from a known environment.
 To establish this threshold, we consider the $q$-th percentile of scores obtained from the validation set.
@@
 3.1 Domain Generalization
-The aim of Domain Generalization (DG) is to train models that generalize well under distribution shifts [Muandet et al., 2013, Zhou et al., 2022].
-In contrast to domain adaptation (DA) [Wang and Deng, 2018], where samples from the test domain are given during training, in DG no knowledge about the test environment is available during training.
-Both, DG and DA, involve access to data from multiple domains during training exploiting the heterogeneity between domains.
-For an overview on DG see for instance Wang et al.
-[2022], Zhou et al.
-[2022].
-As a middle-ground between DA and DG, test-time adaptation (TTA) [Liang et al., 2023] and marginal transfer learning [Blanchard et al., 2021] involves the provision of unlabeled samples during test time.
-In TTA, the model is typically fine-tuned according to the provided samples [Liang et al., 2023].
-Marginal transfer learning in DG refers to the setting where access to the marginal feature distribution is given during testing and provided to the model via $\frac{1}{n}\sum_{i=1}^{n}\sigma(\mathbf{X}_{i})$ [Blanchard et al., 2021].
-In this context, $\sigma$ represents a feature extraction model applied to each $\mathbf{X}_{i}$, and the summation ensures that the operation remains invariant under permutations.
-Different choices for $\sigma$ are explored in the literature.
-Specifically, Zhang et al.
-[2021] utilizes CNNs, Blanchard et al.
-[2021], Dubey et al.
-[2021] incorporates kernel embeddings, and Bao and Karaletsos [2023] adopts patch embeddings.
-While these works highlight the advantages of their methodologies, they fall short in analyzing the exact settings and conditions that contribute to these benefits compared to a standard model.
-Additionally, they do not address the detection of potential failure cases or the utilization of context embedding for model selection.
-A recent method proposes to use transformers to exploit previously seen samples instead of learning permutation-invariant embedding [Gupta et al., 2023]
-The trade-off between ID performance and OOD performance is well-known [Müller et al., 2022, Magliacane et al., 2018, Zhang et al., 2023].
-A method to mitigate this trade-off was proposed in Zhang et al.
-[2023].
-Their approach presupposes knowledge of the data’s originating environment, whereas our goal is to infer it.
+Domain Generalization (DG) trains models to perform under distribution shifts without access to test environments [5, 6].
+In contrast, Domain Adaptation (DA) assumes test samples are available during training [14].
+Both exploit multiple source domains, but DG is strictly test-agnostic.
+Non-marginal DG approaches fall into three groups [15]: data manipulation [16, 17], robust representation learning [18, 19], and learning strategy modification [20, 21] (see 15, 6 for reviews).
+Between DA and DG lie test-time adaptation (TTA) and marginal transfer learning.
+TTA adapts to unlabeled test samples, often via fine-tuning or domain metadata [22, 23].
+Marginal transfer instead assumes access to the marginal feature distribution $\frac{1}{n}\sum_{i}\sigma(\mathbf{X}_{i})$ [7], with $\sigma$ implemented via CNNs [24], kernel embeddings [7, 25], or patch embeddings [26].
+While Blanchard et al.
+[7] analyze kernel embeddings theoretically, existing work leaves open conditions for effectiveness, failure detection, and context-aware model selection.
+A recent alternative replaces permutation-invariant embeddings with transformers that exploit sample order [27].
+Marginal transfer parallels in-context learning: labeled samples define context in the latter [28, 29], while unlabeled samples do so in the former.
+Finally, balancing in-domain and out-of-domain performance remains a central challenge [11, 12, 30].
+Methods like Zhang et al.
+[30] mitigate this trade-off when domain identity is known, whereas our goal is to infer it.
 3.2 Learning Permutation-Invariant Representations
-Analyzing set-structured data with neural networks has received much theoretical [Wagstaff et al., 2022, Bloem-Reddy and Teh, 2020, Murphy et al., 2018] and empirical [Zaheer et al., 2017, Lee et al., 2019, Zare and Van Nguyen, 2021] momentum in recent years.
-For instance, Zare and Van Nguyen [2021] build on the set transformer architecture [Lee et al., 2019] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling.
-Differently, dan Guo et al.
-[2021] proposes to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
-Notably, the methods above impose no probabilistic structure on the set-representations, since the latter are mainly used as deterministic features for downstream tasks.
-In the Bayesian literature, sets represent finitely exchangeable sequences, embodying the core probabilistic structure of most Bayesian models [Orbanz and Roy, 2014].
-Typically, hierarchical (aka multi-level) Bayesian models are used to deal with Simpson’s paradox by organizing observations into clusters or levels [Wikle, 2003, Gelman et al., 2013], mirroring the notions of domain or environment in DG.
-Indeed, hierarchical Bayesian models have been successfully applied in many areas of science, but are typically constrained to linear or generalized linear models (GLMs), prioritizing interpretability over predictive performance.
-The machine learning literature is rich with variational methods for learning invariant summary statistics from set-structured data [Edwards and Storkey, 2016, Garnelo et al., 2018, Kim et al., 2019, 2021, Zeng et al., 2022].
-Crucially, none of the above methods pursues the concrete goal of our work, which is improving domain generalization performance through set-based environment representations.
-To achieve this goal, we employ variants of the DeepSet [Zaheer et al., 2017] and SetTransformer [Lee et al., 2019] architectures for our backbone set-encoder throughout all experiments.
-Interestingly, we observe that our approach is widely robust to the choice of set-encoder architectures in the problems considered.
+Analyzing set-structured data with neural networks has received much theoretical [31, 8, 32] and empirical [33, 34, 35] momentum in recent years.
+For instance, [35] build on the set transformer architecture [34] and augment the attentive encoder with the capability to learn dynamic templates for attention-based pooling.
+Differently, [36] proposes to learn set-specific representations, along with global “prototypes”, using an optimal transport (OT) optimization criterion.
+A set-embedding can also be understood as a learned proxy variable for the confounder $E$.
+Generic proxy variables for confounding variables have been explored in the context of estimating the causal effect from $\mathbf{X}$ to $\mathbf{Y}$ in [37, 38].
+While their work focuses on eliminating the effect of the confounding variable $E$, our objective is to leverage it for prediction purposes.
+Furthermore, they require $\mathbf{X}$ causing $\mathbf{Y}$ which does not conform to all prediction tasks.
+We do not require that $\mathbf{X}$ causes $\mathbf{Y}$ in our theoretical analysis and therefore include more scenarios (e.g., when $\mathbf{Y}$ is causing $\mathbf{X}$).
+3.3 OOD Detection and Selective Classification
+Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [39, 40, 41, 42, 43].
+Flagging OOD instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [40].
+OOD detection is closely related to inference with a reject option (also termed selective classification) [44, 45], which allows classifiers to refrain from predicting ambiguous or novel conditions [46].
+The reject option has been extensively studied in statistical and machine learning [47, 48, 49, 50], with early work dating back to the 1950s [51, 52, 47].
+More recently, [9] explored selective classification in DG settings.
+They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent.
+In this work, we consider a post-hoc score based on the $k$-nearest neighbours to the training set in feature space similar to [53], which applies to both classification and regression settings.
+Unlike the approach taken in [9], where the focus lies on features of individual instances, we consider the set summary provided by the set-encoder.
+Thus, we can identify novel environments even when singleton inputs lack sufficient information.
+4 Experiments
+In the following, we explore various aspects of context-aware models.
+First, we show on two datasets that a context-aware model achieves improved performance in ID and OOD settings compared to a baseline model when the necessary conditions of a source component shift are met.
+Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model.
+We also show that novel environment detection can be utilized to avoid failure cases.
+Third, we demonstrate that the necessary criteria (see Section 2.3) can be validated empirically, identifying cases where no benefits of the method can be expected.
+Experimental details can be found in the Appendix and the source code is available at  [^1] https://github.com/LarsKue/context-aware-domain-generalization.
 Figure 2: Experiment 1.
 Left: Toy dataset that conforms to our theoretical criteria.
-A detailed description can be found in subsection E.1.
-Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3.
-Marginal distributions of the five environments are shown on top.
-Right: Comparison of environment classification accuracy for different set sizes.
+Without environmental information, the marked input at $x=2.5$ could belong to either one of the domains numbered 1, 2, or 3 as indicated by the marginal distributions shown on top.
+Right: Comparison of environment classification accuracy for a baseline model versus a mean-pooled set-encoder using different set sizes.
 Distances between environments refer to the distance between the means of the environments.
-3.3 OOD Detection and Selective Classification
-Detecting unusual inputs that deviate from the examples in the training set has been a long-standing problem of conceptual complexity in machine and statistical learning [Aggarwal and Yu, 2001, Yang et al., 2021, Shen et al., 2021, Han et al., 2022, Yang et al., 2022a].
-Flagging out-of-distribution (OOD) instances involves identifying uncommon data points that might compromise the reliability of machine learning systems [Yang et al., 2021].
-OOD detection is closely related to inference with a reject option (also termed selective classification) [Geifman and El-Yaniv, 2017, El-Yaniv et al., 2010], which allows classifiers to refrain from predicting ambiguous or novel conditions [Hendrickx et al., 2021].
-The reject option has been extensively studied in statistical and machine learning [Hellman, 1970, Fumera and Roli, 2002, Grandvalet et al., 2008, Wegkamp and Yuan, 2012], with early work dating back to the 1950s [Chow, 1957, 1970, Hellman, 1970].
-More recently, Müller et al.
-[2023] explored the utility of selective classification in DG settings.
-They investigated various post-hoc scores to define a “competence region” in feature space where a classifier is deemed competent.
-In this work, we consider a post-hoc score based on the $k$-nearest neighbours to the training set in feature space similar to Sun et al.
-[2022], which applies to both classification and regression settings.
-Unlike the approach taken in Müller et al.
-[2023], where the focus lies on features of individual instances, we consider the set summary provided by the set-encoder.
-Thus, we can identify novel environments even when singleton inputs lack sufficient information.
-4 Experiments
-In the following, we explore various aspects of the context-aware model across three different dimensions.
-First, we show on two datasets that a context-aware model achieves improved performance in the ID as well as the OOD setting compared to a baseline model when the conditions of a source component shift are met.
-Second, we show how novel environments can be detected to select between the most predictive (in the ID setting) and the most robust (in the OOD setting) model.
-We also show that novel environment detection can be utilized to avoid failure cases.
-Third, we demonstrate that the necessary criteria (see subsection 2.3) can be validated empirically, identifying cases where no benefits of the method can be expected.
-Experimental details can be found in the Appendix.
-Furthermore, in Appendix B, we conduct an additional experiment resembling Experiment 1, but with high-dimensional inputs
+Smaller distances produce stronger overlap of the marginals.
+A detailed description can be found in Section E.1.
+4.1 Evaluation Approach
+To approximate Criterion 2.1, Criterion 2.2, and Criterion 2.3, we train five models (Table 1).
+Our context-aware model $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ leverages context sets, while the baseline $f^{\mathbf{Y}\mid\mathbf{X}}$ ignores them.
+Their relative improvement is
+$\mathcal{R}_{\text{I}}=\frac{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})},$ (2)
+where $\mathcal{M}(\cdot)$ is a test performance metric (negative L2-loss for regression).
+$\mathcal{R}_{\text{I}}>0$ indicates that Criterion 2.1 holds.
+For Criterion 2.2, we compare the contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ with its baseline $f^{E\mid\mathbf{X}}$:
+$\mathcal{R}_{\text{II}}=\frac{\mathcal{M}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{M}(f^{E\mid\mathbf{X}})}{\mathcal{M}(f^{E\mid\mathbf{X}})}.$ (3)
+We set $n$ such that $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ achieves nearly perfect ID accuracy; $\mathcal{R}_{\text{II}}>0$ supports Criterion 2.2.
+Finally, to test Criterion 2.3, we introduce the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ and compute
+$\mathcal{R}_{\text{III}}=\frac{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (4)
+These relative improvement metrics serve as proxies for the theoretical criteria: when $\mathcal{M}$ is cross-entropy under optimal models, they align with mutual information measures.
+However, if $\mathcal{M}$ accuracy, then this proxy is not isomorphic to the criterion it approximates.
+Model Symbol Description Purpose
+Context-aware (ours) $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $\mathbf{Y}$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Test Criterion 2.1
+Baseline $f^{\mathbf{Y}\mid\mathbf{X}}$ Predicts $\mathbf{Y}$ from $\mathbf{X}$ Reference
+Environment-oracle $f^{\mathbf{Y}\mid\mathbf{X},E}$ Predicts $\mathbf{Y}$ from $\mathbf{X}$ and $E$ Test Criterion 2.3
+Contextual env. $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $E$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Test Criterion 2.2
+Baseline env. $f^{E\mid\mathbf{X}}$ Predicts $E$ from $\mathbf{X}$ Reference for Criterion 2.2
+Table 1: Five models used to evaluate our approach and verify the theoretical criteria.
 Figure 3: Experiment 1.
-Relative improvement of set-encoder (shown in I) approach versus baseline model (0 means no improvement is achieved) on a toy example.
+Relative improvement of marginal transfer learning (shown in I) versus a baseline model (0 means no improvement is achieved) on a toy example.
 We also show I (OOD) on OOD data.
@@
 III demonstrates the relative improvement in predicting the environment when using contextual information compared to the absence of it.
-Sampling variation arises from using different seeds to partition the ID data into training, test and validation set.
-4.1 Evaluation Approach
-To approximate Criterion 2.1, Criterion 2.2 and Criterion 2.3, we are required to train five distinct models (see Table 3 for an overview).
-We denote our composite model as $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ (see Figure 1) and the baseline model (having no access to the context) as $f^{\mathbf{Y}\mid\mathbf{X}}$.
-Based on these two models, we can compute the relative improvement achieved by our model relative to a baseline model via
-$\mathcal{R}_{\text{I}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}.$ (2)
-Here, $\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}})$ denotes a performance measure for our model and similarly for the baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$, both measured on an unseen test set.
-$\mathcal{R}_{\text{I}}>0$ signifies an advantage attained by the context-aware approach and therefore the fulfillment of Criterion 2.1.
-In the regression setting, we consider the negative L2-Loss as the performance measure.
-To validate Criterion 2.2, we train a contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ utilizing both the set input $\mathbf{S}^{(n)}$ and the target input $\mathbf{X}$ to predict the environment label $E$.
-Additionally, we train a baseline environment model $f^{E\mid\mathbf{X}}$ aimed at predicting $E$ solely from $\mathbf{X}$.
-We then compute the relative improvement $\mathcal{R}_{\text{II}}$ [^2] The exact definition for $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ can be found in Appendix D of the contextual environment predictor relative to the baseline environment predictor.
-$\mathcal{R}_{\text{II}}>0$ indicates that Criterion 2.2 is satisfied.
-In our experiments, we choose the set size $n$ such that we achieve approximately 100% accuracy for our contextual environment predictor $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ on ID data.
-Similarly, we consider an environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ that aims to predict $\mathbf{Y}$ from the singleton input and the environment label $E$.
-We define the relative improvement $\mathcal{R}_{\text{III}}$ of the environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ compared to the baseline method $f^{\mathbf{Y}\mid\mathbf{X}}$.
-In this case, the relative improvement $\mathcal{R}_{\text{III}}$ is associated with Criterion 2.3.
+Sampling variation arises from using different seeds to partition the ID data into training, test, and validation sets.
 4.2 Experiment 1: Toy Example
 Setup
-To set the stage, we consider a dataset shown in Figure 2 and explained in detail in subsection E.1.
+To set the stage, we consider a dataset shown in Figure 2.
 The dataset includes data from five different environments, defined by distinct Gaussian distributions.
@@
 mean vector).
-The dataset exemplifies Simpson’s paradox, wherein a naive fit of the data (without accounting for environmental factors) would yield a negatively sloped line (see Figure 10).
-Conversely, using environmental information for prediction enables to capture the true underlying model with positively sloped lines, counteracting the overall negative trend of the mean vectors.
-Importantly, the dataset meets our necessary criteria: We cannot infer the origin environment from a single input alone, as indicated by the overlap between the marginal distributions of $X$, obtained by projecting the entire sample onto the $x$-axis in Figure 2.
-Thus, the setting aligns with Criterion 2.2, and corresponds to Criterion 2.3, since knowing the environment location $\boldsymbol{\mu}^{(j)}$ should improve prediction.
+The dataset exemplifies Simpson’s paradox, wherein fitting without accounting for environmental factors would yield a negatively sloped line.
+This trend reverses to multiple positively sloped lines when considering environmental factors (see Figure 11).
+Importantly, the dataset meets our necessary criteria, since the environment cannot be inferred from a single input as indicated by the overlap of the marginal distributions in Figure 2.
+The mathematical details underlying this dataset are described in Section E.1.
 Results
@@
 As expected, a decrease in the distance between environment marginal means necessitates more samples to differentiate between environments.
-Interestingly, the particular choice of architecture for the permutation-invariant network does not seem to play a significant role for predicting the environment label well, as demonstrated in Appendix I.
-Figure 4: Experiment 1. Predictions performed on the toy dataset illustrated in Figure 2. We show predictions made by both our set-encoder approach and the vanilla model in the ID and OOD settings.
 Next, we assess the predictive capabilities of the context-aware approach across all possible scenarios of “leave-one-environment-out”.
 This involves training on all environments except one and treating the excluded environment as a novel OOD scenario.
-Here, we consider linear models to ensure an optimal inductive bias for the problem (non-linear models achieve similar results, as shown in subsection E.3).
+Here, we consider linear models to ensure an optimal inductive bias for the problem.
 We can see that Criterion 2.1, Criterion 2.2 and Criterion 2.3 are satisfied in Figure 3.
@@
 This aligns with our expectations, as the set input does not offer more information about the target value than the environment label itself.
-Note that for III we achieve less relative improvement since we consider the accuracy here and not the L2-Loss.
-Finally, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 4 for one trained model.
-Our model captures and utilizes the characteristics of each environment for prediction.
-In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences.
-Note that we obtained the best results by considering a class of linear models that aligns with the data-generating process.
-However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see subsection E.3).
-4.3 Experiment 2: Colored MNIST
-Setup
-The ColoredMNIST dataset [Arjovsky et al., 2019] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (all standard labels $<5$ are assigned to new label 0, and all labels $\geq 5$ are the new label 1).
-Furthermore, label noise is deliberately added, so only in 75% of all cases, the label can correctly be predicted from the shape.
-To make things more challenging, the image background can take two colors that are also associated with the image label.
-In the first environment, the association is 90% and in the second one 80%.
-Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape.
-However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy – worse than random.
-This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [Arjovsky et al., 2019, Zhang et al., 2023].
-For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of diminished accuracy in the first two environments (75% instead of 80% or 90%).
-In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
-[…]
+Note that for metric III, we achieve less relative improvement since we consider the accuracy and not the L2-Loss.
+In Appendix E all scenarios where one environment is left out for testing can be found.
+Additionally, we present there similar results for non-linear models and also demonstrate that the specific choice of permutation-invariant network does not significantly impact the prediction of the environment label.
+Furthermore, in Appendix B, we conduct an additional experiment resembling Experiment 1, but with high-dimensional inputs and achieve similar results.
+Accuracy $\%$] $\uparrow$
 ID OOD
@@
 Selection (Baseline) $\mathbf{84.0}\pm 0.3$ $14.0\pm 0.4$
-Bayes Optimal Classifier $\mathbf{85.0}$ $\mathbf{75.0}$
+Bayes Optimal $85.0$ $75.0$
 Table 2: Experiment 2.
-Mean and standard deviation of accuracy percentages across model types and domain settings over 5 runs.
-The features extracted by our model allow for improved OOD detection compared to the features of the baseline model.
-Thus, our model can perform a favorable selection between the baseline model in the ID setting and the invariant model in the OOD setting.
+Accuracy across model types and domain settings.
+Our context-aware model yields improved OOD detection compared to the baseline, allowing model selection at inference time.
+See Appendix K for more details.
+4.3 Experiment 2: Colored MNIST
+Setup
+The ColoredMNIST dataset [54] is an extension of the standard MNIST dataset, wherein the number of classes is reduced to two classes (digits <5<5 and $\geq 5$).
+Furthermore, label noise is deliberately added, such that only in 75% of all cases, the label can be correctly predicted from the shape.
+To make things more challenging, the image background can take two colors that are also associated with the image label.
+In the first environment, the association is 90% and in the second one 80%.
+Therefore, a baseline model would tend to utilize the background for prediction instead of the actual shape.
+However, in a third environment, the associations are reversed, so that a model based on the background color would achieve only 10% accuracy (i.e., worse than random).
+This dataset implies a trade-off between predictive performance in ID domains versus robustness in OOD domains, as discussed in [54, 30].
+For instance, an invariant model that relies solely on an object’s shape would be robust to domain shift at the cost of lower accuracy in the first two environments (75% vs.
+80% or 90%).
+In contrast, a baseline model would achieve greater accuracy in the first domains (80% and 90%), but would fail dramatically in the third domain (only 10%).
 Results
-Here, we assume the invariant model to be given (see Appendix G for details), but it could also be obtained by invariant learning, e.g.
-Invariant Risk Minimization [Arjovsky et al., 2019].
-With our novel environment detection approach (see subsection 2.5) we can get the best of both worlds, circumventing the inherent trade-off.
+Here, we assume the invariant model to be given, but it could also be obtained by invariant learning, e.g.
+Invariant Risk Minimization [54].
+With our novel environment detection approach (see Section 2.5) we can get the best of both worlds, circumventing the inherent trade-off.
 When identifying the ID setting, we utilize the baseline model that achieves the highest predictiveness within the observed environments.
@@
 The results can be found in Table 2.
-By utilizing the model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data.
+By utilizing model selection based on the set-summary $h_{\boldsymbol{\psi}}(\mathbf{S}^{(n)})$, we nearly recover the ID accuracy while maintaining identical performance to the invariant model on OOD data.
 Evidently, the novel environment detection only works with set summaries.
 A feature extracted from a single sample does not provide enough information to reliably detect distribution shifts, leading to difficulties in effectively selecting between baseline and invariant model, as demonstrated in Table 2.
+Details on this experiment can be found in Appendix G.
 4.4 Experiment 3: Violated Criteria
 Setup
-We consider the PACS dataset [Li et al., 2017], training our model on the Cartoons, Sketches, and Paintings environments, and assess its performance in the Art environment during testing.
-The dataset includes images with labels that we intend to predict.
-Regarding a second classification task, we delve into the OfficeHome dataset [Venkateswara et al., 2017].
-In line with the PACS dataset, we approach the classification problem, training across three specific environments, and subsequently evaluating a novel one as an out-of-distribution (OOD) scenario.
+To demonstrate the effects of criterion violation, we consider the PACS dataset [55], as well as the OfficeHome dataset [56], each with the Art environment chosen as the novel (OOD) domain.
+Results
+As expected, when the criteria are not met, context-aware models cannot achieve a benefit over the baseline (see Figure 4).
+Validating the criteria empirically, we find that Criterion 2.2 is not satisfied for PACS, as a single sample is sufficient to infer the source domain with near-perfect accuracy.
+Furthermore, Criterion 2.3 is not satisfied, as 4(a) depicts.
 (a) Environment Art in PACS dataset.
@@
 (b) Environment Product in OfficeHome dataset. Although the environment is not inferable from one input sample (Criterion 2.2), the environment information does not yield benefits (Criterion 2.3).
-Figure 5: Experiment 3.
-Tell-tale examples where at least one of the necessary criteria is not satisfied and the context-aware approach cannot possibly yield benefits.
-For experimental details, see Appendix H.
-Results
-When the criteria are not met, no benefits can be achieved, even in the ID setting.
-This has been proven in Theorem 2.1 and we demonstrate it here for two scenarios empirically (see Figure 5).
-We find that Criterion 2.2 is not satisfied on the PACS dataset: As depicted in 4(a), the contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ does not perform better compared to the baseline environment model $f^{E\mid\mathbf{X}}$.
-Remarkably, a single example is sufficient to infer the source environment, allowing for a 99.7% accuracy in predicting the correct environment from an individual sample (see Appendix H).
-Also Criterion 2.3 is not satisfied as 4(a) depicts.
-Since the criteria are not met, we do not achieve any benefit over the baseline model, neither in the ID nor in the OOD setting, as demonstrated in 4(a).
+Figure 4: Experiment 3. Tell-tale examples where at least one of the necessary criteria is not satisfied and the context-aware approach cannot possibly yield benefits.
 On the OfficeHome dataset we find that Criterion 2.2 is not satisfied, while Criterion 2.3 is.
 Results are depicted in 4(b).
-We observe that the set input offers benefits for predicting the data originating environment corresponding to Criterion 2.3.
-However, even when providing the target classifier with the environment label (environment-oracle model), we do not achieve an improvement over the baseline model suggesting that Criterion 2.2 is not satisfied.
-As expected and depicted in 4(b), our method does not yield benefits compared to the baseline model.
+We observe that the set input offers benefits for predicting the source environment corresponding to Criterion 2.3.
+However, even when providing the target classifier with the environment label, we do not achieve a benefit, suggesting that Criterion 2.2 is not satisfied.
+For experimental details, see Appendix H.
 4.5 Experiment 4: Failure Case Detection
+Winter MSE $\downarrow$ AUROC $\%$] $\uparrow$
+ID OOD
+Baseline $2.21\pm 0.11$ $6.08\pm 0.13$ $58.2\pm 0.7$
+Ours $\mathbf{2.09}\pm 0.12$ $\mathbf{5.7}\pm 0.4$ $\mathbf{100.0}\pm 0.0$
+Table 3: Experiment 4.
+Inference performance (MSE) and novel environment detection (AUROC) comparison between our context-aware model and the baseline for the winter domain in the BikeSharing dataset.
+See Appendix K for more details.
 Setup
@@
 This might be unattainable by the model, making the inclusion of a reject option beneficial.
-Using the BikeSharing dataset [Fanaee-T, 2013], we demonstrate that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary.
-For this dataset we consider the task of predicting the number of bikes rented across the day based on weather data.
+Using the BikeSharing dataset [57], we demonstrate that in cases where different seasons like summer or winter represent distinct environments, extrapolation might be necessary.
+We consider the task of predicting the number of bikes rented across the day based on weather data.
 Here we explore the scenario where we train on all seasons except winter.
-Details about the dataset, pre-processing steps and other scenarios can be found in Appendix J.
+Details about the dataset, pre-processing steps, and other testing scenarios can be found in Appendix J.
 Results
@@
 However, both the baseline and the context-aware approach experience performance degradation in the novel winter environment.
-To detect the novel environment and, consequentially, potential failure cases, we compute the score as suggested in subsection 2.5 and evaluate how well it distinguishes between ID versus OOD environments.
+To detect the novel environment and, consequentially, potential failure cases, we compute the score as suggested in Section 2.5 and evaluate how well it distinguishes between ID versus OOD environments.
 We designate an independent ID test set and use the environment excluded during training (here winter) as the OOD set for evaluation.
-The area under the ROC-curve (AUROC) in Table 5 demonstrates that the score based on the permutation-invariant embedding allows for a perfect detection of the novel environment, whereas the standard approach fails as expected.
-Table 2: Experiment 4.
-Performance comparison between our model and the baseline for the winter domain in the BikeSharing dataset.
-We compare the performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC).
-Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case.
-We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
-Other domains can be found in Appendix J.
+The area under the ROC-curve (AUROC) in Table 3 demonstrates that the score based on the permutation-invariant embedding allows for perfect detection of the novel environment, whereas the standard approach fails as expected.
 5 Conclusions
 In this work, we aimed to advance the theoretical understanding of marginal transfer learning in domain generalization.
-Thus, we formulated criteria that are necessary for the approach to yield benefits and are also easy to verify empirically.
-Moreover, we pinpoint the source component shift as a scenario where context-aware models can offer advantages, enabling the identification of favorable scenarios.
-Additionally, we illustrate the capability to identify novel environments, thereby allowing for the detection of potential failure cases.
-This, in turn, also supports model selection by choosing the most predictive on ID data and the most robust model on OOD data.
-Acknowledgments
-JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation. We thank Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
-Appendix A Algorithm
+Accordingly, we formalized criteria that are necessary for context-aware models to yield benefits and are also verifiable in practice.
+Moreover, we pinpointed the source component shift as a scenario where context-aware models can offer advantages, enabling the identification of favorable settings and the identification of potential failure cases.
+The latter allows us to perform real-time model selection between the best performing model on ID data and the most robust (i.e., domain-invariant) model on OOD data.
+Future research should investigate generalization bounds, the learner’s behavior in finite-data regimes, and the generalization behavior of the learner as the number of training domains increases (i.e., the domain efficieincy).
+Acknowledgments and Disclosure of Funding
+This work is partially funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project 528702768.
+JM and UK were supported by Informatics for Life funded by the Klaus Tschira Foundation.
+We thank Paul Christian Bürkner, Florian Fallenbüchel, Felix Draxler, and Armand Rousselot for their support and fruitful discussions.
+Appendix A Pseudocode
 Data: Samples from the joint distribution $p(\mathbf{X},\mathbf{Y},E)$
 Input : Composite model parameters $\boldsymbol{\theta}$, set size $n$, batch size $m$, loss-function $c$, number of iterations $k$, learning rate schedule $\alpha(k)$
-for $i=1,\dots,k$ do
-       Sample mini-batch =$\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
-       for $j=1,\dots,m$ do
-             Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
-             Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
-       end for
-      Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
+1 for $i=1,\dots,k$ do
+2   Sample mini-batch $\mathcal{B}=\{(\mathbf{x}_{1},\boldsymbol{y}_{1},\text{env}_{1}),\dots,(\mathbf{x}_{m},\boldsymbol{y}_{m},\text{env}_{m})\}$ from $p(\mathbf{X},\mathbf{Y},E)$
+3 for $j=1,\dots,m$ do
+4      Sample set $\mathbf{s}_{j}^{(n)}=\{\mathbf{x}_{1},\dots\mathbf{x}_{n}\}$ from $p(\mathbf{X}\mid E=\text{env}_{j})$
+5      Replace $\text{env}_{j}$ with $\mathbf{s}_{j}^{(n)}$ in $\mathcal{B}$
+6   end for
+7  Update $\boldsymbol{\theta}$ using adaptive mini-batch gradient descent (or any variant):
 $\boldsymbol{\theta}_{k}\leftarrow\boldsymbol{\theta}_{k-1}-\alpha(k)\nabla_{\boldsymbol{\theta}}\left(\sum_{j=1}^{m}c\left(f_{\boldsymbol{\theta}}(\mathbf{x}_{j},\mathbf{s}_{j}^{(n)}),\boldsymbol{y}_{j}\right)\right)$
-end for
+8 end for
 Output : Trained context-aware model $f_{\boldsymbol{\theta}}$
 Algorithm 1 Optimizing Equation 1 for context-aware domain generalization.
-A.1 Model Overview
-For a detailed overview of the models used to verify the criteria, see Table 3.
-Model Symbol Description Purpose
-Context-aware model (ours) $f^{\mathbf{Y}\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts Y from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Improve predictions
-Baseline model $f^{\mathbf{Y}\mid\mathbf{X}}$ Predicts Y from $\mathbf{X}$ alone Verify improvement
-Environment-oracle model $f^{\mathbf{Y}\mid\mathbf{X},E}$ Predicts Y from $\mathbf{X}$ and $E$ Verify Criterion 2.3
-Contextual environment model $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ Predicts $E$ from $\mathbf{X}$ and $\mathbf{S}^{(n)}$ Verify Criterion 2.2
-Baseline environment model $f^{E\mid\mathbf{X}}$ Predicts $E$ from $\mathbf{X}$ alone Verify Criterion 2.2.
-Table 3: The five different model types used to evaluate our approach and verify the theoretical criteria for improvement.
 Appendix B Additional Experiment: ProDAS
 Setup
-We utilize the ProDAS library [Müller et al., 2023] to generate high-dimensional image data that meets our dataset requirements.
+We utilize the ProDAS library [58] to generate high-dimensional image data that meets our dataset requirements.
 The dataset comprises objects of shape square and circle, exhibiting variations in their texture, background color, rotation, and size.
@@
 Then, we achieve $\mathbf{Y}\not\perp\mathbf{S}^{(n)}\mid\mathbf{X}$, recovering Criterion 2.1.
-The proof can be found in subsection C.3.
+The proof can be found in Section C.3.
 C.2 Insufficiency of Criteria 2 and 3 for Criterion 1
@@
 Due to the overlap between the third and the first two environments, a set input provides additional information about $E$ compared to a single sample $X$, verifying Criterion 2.2.
-Regarding the mechanism relating inputs to outputs, we assume that on $[a,\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ differs, e.g., two constant functions with distinct values.
+Regarding the mechanism relating inputs to outputs, we assume that on $a,\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ differs, e.g., two constant functions with distinct values.
 We further assume that on $(\frac{a+b}{2},b+\frac{a+b}{2}]$ the relation between input $X$ and output $Y$ does not vary with the environment, e.g., is constant.
-This aligns with Criterion 2.3: if we know the environment, we can improve the prediction, specifically on $[a,\frac{a+b}{2}]$.
+This aligns with Criterion 2.3: if we know the environment, we can improve the prediction, specifically on $a,\frac{a+b}{2}]$.
 However, Criterion 2.1 is not satisfiable.
@@
 Since the relation between $X$ and output $Y$ differs only in the supports of environment 1 and environment 2 (specifically, it differs in $\mathcal{U}[a,\frac{a+b}{2}]$), the set input cannot provide additional information about the output $Y$ compared to the single input $X$, i.e.
-it holds $Y\perp\mathbf{S}^{(n)}\mid X$,
-C.3 Proof of Theorem 2.1
+it holds $Y\perp\mathbf{S}^{(n)}\mid X$.
+It is also worth noting that Criterion 2.3 might be achievable while Criterion 2.2 is unattainable and vice versa.
+For instance, when we can infer the originating environment from one sample (Criterion 2.2 is not attainable), the relation between $\mathbf{X}$ and $\mathbf{Y}$ might still vary with the environment (Criterion 2.3 is achievable).
+C.3 Illustration and Proof of Theorem 2.1
+Figure 6: Illustration of Theorem 2.1.
+The first row depicts (a), the second row (b) and the third row (c).
+The pink framed plots show the conditional distributions along the pink marker as shown on the right.
 In the following, we give proofs of Theorem 2.1 (a) - (d).
@@
 $I(A+B;C)$ $=h(A+B)-h(A+B\mid C)$
-=$\overset{(\star)}{=}\left(h(A)+h(B)-h(A\mid A+B)\right)-\left(h(A\mid C)+h(B\mid C)-h(A\mid A+B,C)\right)$
+$\overset{(\star)}{=}\left(h(A)+h(B)-h(A\mid A+B)\right)-\left(h(A\mid C)+h(B\mid C)-h(A\mid A+B,C)\right)$
 $=I(A;C)+I(B;C)-h(A\mid A+B)+h(A\mid A+B,C)$
 $\overset{(\star\star)}{\leq}I(A;C)+I(B;C)$ (15)
-()($\star$)( ⋆ ) follows with the chain rule for entropy
+$(\star)$ follows with the chain rule for entropy
 $h(A,A+B)$ $=h(A)+h(A+B\mid A)$ (16)
@@
 which implies $h(A+B)=h(A)+h(B)-h(A\mid A+B)$ and equally when conditioning on $C$.
-[…]
-subsection C.3 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
+$(\star\star)$ follows since $h(A\mid A+B,C)\leq h(A\mid A+B)$.
+Section C.3 can be extended to the conditional mutual information if $A\perp B\mid D$ and $A\perp B\mid D,C$:
 $I(A+B;C\mid D)\leq I(A;C\mid D)+I(B;C\mid D)$ (19)
@@
 For one, we require $\mathbf{S}^{(n)}\mid Z\mid\mathbf{X}$.
-This can be interpreted as “increasing the set size does not improve the prediction of $E$”.
+This can be interpreted as “increasing the set size does not improve the prediction of EE”.
 Also $\mathbf{S}^{(n)}\perp Z\mid\mathbf{X},Y$ can be interpreted similarly: increasing the set size and considering the ground truth label/value does not enhance the predictability of $E$.
@@
 We define the relative improvements $\mathcal{R}_{\text{II}}$ and $\mathcal{R}_{\text{III}}$ as
-$\mathcal{R}_{\text{II}}=\frac{\mathcal{L}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{L}(f^{E\mid\mathbf{X}})}{\mathcal{L}(f^{E\mid\mathbf{X}})}$ (22)
+$\mathcal{R}_{\text{II}}=\frac{\mathcal{M}(f^{E\mid\mathbf{X},\mathbf{S}^{(n)}})-\mathcal{M}(f^{E\mid\mathbf{X}})}{\mathcal{M}(f^{E\mid\mathbf{X}})}$ (24)
 and
-$\mathcal{R}_{\text{III}}=\frac{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{L}(f^{\mathbf{Y}\mid\mathbf{X}})}$ (23)
+$\mathcal{R}_{\text{III}}=\frac{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X},E})-\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}{\mathcal{M}(f^{\mathbf{Y}\mid\mathbf{X}})}$ (25)
 $\mathcal{R}_{\text{II}}$ signifies the relative performance gain in predicting the environment when the set input is given compared to the solitude input.
@@
 Here we only consider the datasets where we performed a full evaluation of all criteria.
+D.1 Computational complexity
+We run all experiments using four Titan X GPUs, with 12GB VRAM each.
+On this hardware, each experiment requires between two and three days to run to completion.
+Our code base provides several utilities to reduce the overall memory footprint, allowing reproduction of our experiments on less powerful hardware.
 Appendix E Experiment 1: Details
 E.1 Data Generation
-Simpson’s Paradox [Peters et al., 2017, von Kügelgen et al., 2021] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined.
-There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [Bickel et al., 1975] or an evaluation of the efficacy of different treatments for kidney stones [Charig et al., 1986].
-To replicate this, we create a dataset inspired by an illustration of Simpson’s Paradox on Wikipedia sim .
+Simpson’s Paradox [3, 4] describes a statistical phenomenon wherein several groups of data exhibit a trend, which reverses when the groups are combined.
+There are several famous real-world examples of Simpson’s Paradox, such as a study examining a gender bias in the admission process of UC Berkeley [59] or an evaluation of the efficacy of different treatments for kidney stones [60].
+To replicate this, we create a dataset inspired by an illustration of Simpson’s Paradox on Wikipedia [61].
 The dataset consists of a mixture of 2D multivariate normal distributions, with the intent of using the first dimension as a feature, and the second as a regression target.
@@
 Setting Value Controls
-n_domains $5$ number of mixture components
+n_domains 55 number of mixture components
 n_samples $10000$ number of samples per mixture component
@@
 Table 4: Default Settings for the Simpson’s Paradox Dataset. Samples from the dataset constructed with these settings can be seen in Figure 2
-The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [minutephysics and Reich, 2017].
+The YouTube channel minutephysics also published a short descriptive video on this phenomenon in 2017 [62].
 E.2 Training Details
@@
 We choose the mean as the pooling operation.
+Figure 7: Experiment 1. Predictions performed on the toy dataset illustrated in Figure 2. We show predictions made by both our set-encoder approach and the vanilla model in the ID and OOD settings.
+Now, we visualize the predictions of the baseline approach and our set-encoder approach in Figure 7 for one trained model.
+Our model captures and utilizes the characteristics of each environment for prediction.
+In contrast, the baseline approach struggles to discern between environments due to the significant overlap between environments, resulting in an inability to deal with environmental differences.
+Note that we obtained the best results by considering a class of linear models that aligns with the data-generating process.
+However, we observe that extrapolation performance drops when the considered models are overly complex and lack a strong inductive bias (see Section E.3).
 E.3 Non-Linear Models
-In the experiments in subsection 4.2, we considered linear models for our model and the baseline.
+In the experiments in Section 4.2, we considered linear models for our model and the baseline.
 In the following, we show results for the non-linear model class in Figure 8.
@@
 To select between the baseline model and the invariant model, we are required to distinguish between ID and OOD data.
-Therefore, we follow the approach proposed in subsection 2.5.
+Therefore, we follow the approach proposed in Section 2.5.
 We consider the $k$-nearest neighbors of the training set to compute the score $s_{\boldsymbol{\psi}}$ where $k=5$.
 Since we compare the scores elicited by features of the baseline model with the scores elicited by the features extracted by the set-encoder, we restricted both architectures to have the same feature dimension.
-To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see subsection 2.5 for details).
+To establish a threshold for distinguishing between ID and OOD samples, we designate samples with scores below the 95% quantile of the validation set as ID and those above as OOD (see Section 2.5 for details).
 In total, we explore five dataset seeds to partition into training, validation, and test sets.
@@
 For the PACS as well as the OfficeHome dataset we consider features $g(\mathbf{X})$ that are kept fixed and not optimized.
-Here, we employ the Clip features proposed in Radford et al.
-[2021].
+Here, we employ the Clip features proposed in [63].
 The inference model, baseline model, and set-encoder are kept simple and employ only linear layers followed by ReLU activation functions.
@@
 We compare their performance in the ID and OOD setting (MSE), as well as their capability to detect a novel environment (AUROC).
-Both models fail in the OOD setting, but our model can detect with strong certainty when this is the case.
-We present the mean and standard deviation derived from 5 runs using different seeds for partitioning into training, validation, and test sets.
+Both models experience a performance drop in the OOD setting, but our model can detect with strong certainty when this is the case.
+See Appendix K for more details.
 Dataset / Set Size Simpson / $32$
@@
 $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$ $\mathbf{100.0}\pm 0.0$
-Dataset / Set Size ProDAS / $128$ OfficeHome / $4$ PACS / $4$
+Dataset / Set Size ProDAS / $128$ OfficeHome / 44 PACS / 44
 Domain 1 2 3 4 Product Art
-$f^{E\mid\mathbf{X}}$ $43.8\pm 1.1$ $50.0\pm 1.3$ $49.9\pm 2.3$ $44.4\pm 1.0$ $86.16\pm 0.33$ $\mathbf{99.72}\pm 0.33$
+$f^{E\mid\mathbf{X}}$ $43.8\pm 1.1$ $50.0\pm 1.3$ $9.9\pm 2.3 44$$4.4\pm 1.0$ $86.16\pm 0.33$ $\mathbf{99.72}\pm 0.33$
 $f^{E\mid\mathbf{X},\mathbf{S}^{(n)}}$ $\mathbf{99.6}\pm 0.6$ $\mathbf{99.5}\pm 1.0$ $\mathbf{98.7}\pm 1.6$ $\mathbf{98.0}\pm 3.2$ $\mathbf{98.49}\pm 0.24$ $\mathbf{100.0}\pm 0.0$
@@
 Domains that do not overlap exactly will remain distinguishable, so long as the set size is large enough.
-Therefore, we construct the toy dataset as described in subsection E.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
+Therefore, we construct the toy dataset as described in Section E.1, but use the setting n_domains = 2 and vary the distance between environments for each experiment.
 We train each architecture on this dataset for $20$ epochs, using $5$ different seeds.
@@
 We choose the batch size at $128$ fixed.
-Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [Agarap, 2019] activation.
+Each architecture consists of a linear projection into a $64$-dimensional feature space, followed by a fully connected network with $3$ hidden layers, each containing $64$ neurons and a ReLU [64] activation.
 For the set-based methods, this is followed by the respective pooling.
 We choose $8$ heads for the attention-based model.
-Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [Zhang and Sabuncu, 2018].
+Finally, the output is linearly projected back into the $2$-dimensional logit space, where the loss is computed via cross-entropy [65].
 For methods that support a non-unit output set size, we choose the output set size as $4$. The output set is mean-pooled prior to projection into the logit space.
@@
 Appendix J Bike Sharing Dataset
-This dataset, taken from the UCI machine learning repository [Fanaee-T, 2013], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bikeshare system.
+This dataset, taken from the UCI machine learning repository [57], consists of over $17000$ hourly and daily counts of bike rentals between 2011 and 2012 within the Capital bike share system.
 Each dataset entry contains information about the season, time, and weather at the time of rental. Casual renters are also distinguished from registered ones.
-Similar to Rothenhäusler et al.
-[2021], we only consider the hourly rental data.
+Similar to [66], we only consider the hourly rental data.
 We drop information about the concrete date and information about casual versus registered renters.
 We choose the season variable (spring, summer, fall, winter) as the environment and the bike rental count as the regression target.
-Since we deal with count data, we also apply square root transformation on the target similar to Rothenhäusler et al.
-[2021].
+Since we deal with count data, we also apply square root transformation on the target similar to [66].
+Appendix K Table Details
+For tables 2, 3, and 5, we present the mean and standard deviation computed over 5 different training runs using separate seeds for partitioning the data into training, validation, and test sets.
+We compute the AUROC by calculating a score for each sample as described in Section 2.5.
+The AUROC is then determined by calculating the AUC of the ROC curve, which is associated with the task of predicting the environment.
+We highlight models within the 95% confidence interval of the best one for each respective category in bold.
+Appendix L Potential Societal Impacts
+This paper presents a foundational study, with societal impacts reliant mostly on the application of our methods.
+Nevertheless, we estimate that good-faith applications of our methods can have a positive societal impact.
+This manifests in improved performance results when our criteria are satisfied, as well as increased trustworthiness of these results due to the reliant detection of novel environments.
+This is particularly important for safety-critical applications, e.g., in medicine.
+Negative societal impacts may also manifest in bad-faith applications, as the improved results may be misused.
+Furthermore, there is a risk that our methods may inadvertently perpetuate existing biases in data, particularly if environments are chosen in bad faith.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: We claim to empirically and theoretically analyze the conditions under which set-encodings can benefit marginal transfer learning.
+We show this via mathematical proofs and on a range of experiments, including possible failure cases.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: We provide an extensive discussion of the limitations of our approach throughout the paper.
+For instance, we consider the scenario when our theoretical criteria are violated in Section 4.4.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We jointly show our assumptions and proofs in Section C.3.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We fully discuss experimental details, including a description of architectures and parameters, in Appendix D.
+All datasets used are publicly available, and ready to use from within our code base, where we also provide further instructions for reproducibility.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: We provide open access to the experiment code, anonymized for review purposes.
+All datasets used are publicly available, and ready to use from our code base.
+We also provide further instructions to reproduce the experiments in Appendix D and in our code repository.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: We provide extensive detail on experimental settings and parameters in Appendix D.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: For our evaluation metrics, we derive a mean and standard deviation from multiple runs using separate seeds for data splitting.
+When models are within the 95% confidence interval from the best one, we choose to also highlight it in bold, as discussed in Appendix K.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We discuss compute resources in Section D.1.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We carefully reviewed the NeurIPS Code of Ethics and found no ethical concerns for this paper. We discuss potential harmful societal impacts in Appendix L.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discuss potential positive and negative societal impacts in Appendix L.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54.
+Justification: We do not consider our contributions to pose a high risk, as we do not release large-scale models, image generators, or datasets.
+Potential misuse of our methodologies is discussed in Appendix L.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59.
+Justification: We appropriately cite all original papers of methods, datasets, model architectures, evaluation metrics, code repositories, etc.
+Package versions will be made public alongside our code repository upon acceptance.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We do not release new assets with this paper.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing, nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing, nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.